### v1

# Nathalie Gimenes PUC-Rio Tonghui Qi

Abstract
How sellers choose reserve prices is central to auction theory, and the optimal reserve price depends on the seller’s risk attitude. Numerous studies have found that observed reserve prices lie below the optimal level implied by risk-neutral sellers, while the theoretical literature suggests that risk-averse sellers can rationalize these empirical findings. In this paper, we develop an econometric model of ascending auctions with a risk-averse seller under independent private values. We provide primitive conditions for the identification of the Arrow–Pratt measures of risk aversion and an estimator for these measures that is consistent and converges in distribution to a normal distribution at the parametric rate under standard regularity conditions. A Monte Carlo study demonstrates good finite-sample performance of the estimator, and we illustrate the approach using data from foreclosure real estate auctions in São Paulo.
JEL Classification Numbers: C14, C21, C57
Keywords: Ascending Auction, Identification, Local Polynomial Estimation, Quantile Regression, Semiparametric Estimation
1 Introduction
The reserve price is one of the seller’s primary tool for influencing expected revenue in an auction. Under the assumption of a risk-neutral seller, reserve prices play an essential role in the celebrated Revenue Equivalence Theorem of Myerson, (1981), while Riley and Samuelson, (1981) demonstrated that the optimal reserve is strictly positive, even with strong competition among bidders, across a range of auction formats. These insights form the benchmark that has guided much of the subsequent theoretical and empirical analysis of reserve-price behavior.
A number of empirical studies, however, have documented that observed reserve prices are well below the levels implied by the risk-neutral benchmark in a variety of settings. Examples include McAfee and Vincent, (1992), Paarsch, (1997), McAfee et al., (2002), Haile and Tamer, (2003), and Tang, (2011). One explanation is that sellers themselves are risk-averse, leading them to set lower reserves in order to increase the probability of a sale (Maskin and Riley, (1984), Matthews, (1987), Hu et al., (2010)).
Although the econometric literature on auctions has increasingly incorporated risk aversion into structural models, to date, the focus has been on the bidders in the context of first-price auctions (e.g., Lu and Perrigne, (2008), Guerre et al., (2009), Campo et al., (2011), Li et al., (2015), Zincenko, (2018), Grundl and Zhu, (2019)). We are not aware of any prior work that estimates the risk aversion of the seller despite its relevance in the theoretical literature.
In this paper, we propose an ascending auction model with risk-averse seller under the independent private value (IPV) framework. We provide primitive conditions, which parallel those assumed in the theoretical auction literature, under which Arrow-Pratt measures of risk aversion can be identified from the distribution of winning bid, reserve price, and seller’s valuation of the sale object. Subsequently, we show how the seller’s risk-aversion parameter in a parametric model, such as CARA or CRRA utility function, can be consistently estimated at the parametric rate with a limiting normal distribution. Our estimator is a two-step semiparametric estimator, as the risk parameter is identified by the first-order condition the seller uses for setting an optimal reserve, which involves the quantile of the bidder’s valuation that we flexibly estimate in the first step.
We perform a Monte Carlo study and find that our estimator behaves as expected from the theory. We then apply our methodology to real estate foreclosure auction data in São Paulo as an illustration. In this application, the sellers are lenders (banks and private companies) and the Court of Justice of the State of São Paulo. Our finding indicates that sellers are risk-averse, and the model with a risk-averse seller fits the data better than one with a risk-neutral seller.
Beyond explaining low reserve prices, knowing the seller’s risk attitude has policy implications. Auction design matters more when the seller is risk averse. For examples, the Revenue Equivalence Theorem does not hold with a risk-averse seller, and while some optimal auction results in Myerson, (1981) extend to risk-averse sellers, they may do so only in some cases (Sundararajan and Yan, (2020)). The reason for why sellers may be risk-averse depends on applications. In ours, it is straightforward to justify why sellers are averse to not being able to sell foreclosed properties. Firstly, there may be direct costs for holding on to such properties (e.g., property taxes, insurance, security). The value of foreclosure homes also tend to devalue over time. Moreover, from the accounting perspective, foreclosure assets are typically viewed as non-performing loans and sellers may have directives to sell them within a certain time frame [^1] This is indeed a guideline of some central banks, for example, see this report by the ECB https://www.bankingsupervision.europa.eu/ecb/pub/pdf/guidance_on_npl.en.pdf.
There are two steps to estimating our model. The first is the estimation of the bidders’ valuations distribution from bids data. The second is on the estimation of the risk parameter. The first stage of our analysis builds on the large literature on the identification and estimation of bidders’ valuations distribution from bids data, which starts with the parametric model by Paarsch, (1992) and grows with Donald and Paarsch (1993,1996), Laffont et al., (1995), Athey and Levin, (2001), Hirano and Porter, (2003), Li and Zheng, (2012) amongst others. The parametric models, however, can be computationally demanding to estimate and is subjected to model misspecification. The nonparametric approach, which begins with the seminal work of Guerre et al., (2000), offers an alternative that have been developed in a number of contexts, inter alia, see Athey and Haile, (2002), Lu and Perrigne, (2008), Krasnokutskaya, (2011), Marmer and Shneyerov, (2012), Hubbard et al., (2012), Campo et al., (2011), Marmer et al., (2013), Enache and Florens, (2017), Liu and Luo, (2017), Luo and Wan, (2018), Zincenko, (2018) and Ma et al., (2019). The purely nonparametric approach is also not without issues, as it suffers from the curse-of-dimensionality. Given the well-known trade-offs between parametric and nonparametric modelling, our work adopts the augmented quantile regression (AQR) approach of Gimenes and Guerre, (2022) (hereafter GG22) that uses a linear quantile specification to model the distribution of the bidder’s valuation. [^2] GG22 also proposed a more flexible version of the AQR where linear quantile is replaced by nonparametric additive functions.
AQR is a novel way to estimate quantiles by minimizing an integrated version of the classic Koenker and Bassett, (1978) objective function, combined with kernel smoothing and polynomial approximation of the quantile function (cf. Fan (1992,1993)). The local polynomial formulation is appealing to our application as it improves boundary properties and simultaneously delivers estimates of the quantile function and its derivatives; the latters enter the seller’s first-order condition that we use to estimate the seller’s risk aversion parameter. Our AQR estimator of bidder valuation quantiles in ascending auctions in fact targets the same object of interest as in Gimenes, (2017), who estimated it using the classic Koenker and Bassett objective function without local-polynomial features. It should be emphasized that our work does not contribute to the methodological development of the AQR estimator itself. However, our paper goes some way in expanding the range of its applications beyond the results stated in GG22, as the authors’ focus was on estimating bidder valuation quantiles from first-price bids.
In particular, we provide pointwise properties of the AQR quantile estimator and its inverse. The former was not given in GG22, although they provided the uniform convergence rate of the quantile estimator. They did not consider the latter, which serves as a flexible estimator of a conditional CDF and is of independent interest. We present these results in general form and apply them to an ascending auction model. Moreover, our analysis of the seller’s risk-aversion measure, which had not been estimated previously, makes use of Theorem 4 in GG22 that establishes asymptotic normality for integral functionals of AQR estimators. GG22 motivate the study of such functionals with examples that have closed-form expressions in the quantile functions (expected revenue (Li et al., (2003)) and bidder’s risk aversion parameter in a first-price auction (Guerre et al., (2009))). We show that such functionals can also be obtained when the AQR estimator is used to estimate nuisance functions in a general two-step estimation problem.
To estimate the seller’s risk-aversion measure in the second step, we assume that the seller’s utility function is known parametrically. Our risk aversion estimator satisfies a first-order condition that depends on the derivative and inverse of the bidder valuation quantile function, which we estimate in the first stage using the AQR method. To derive its limiting distribution, we take the pathwise differentiation approach to capture the contribution of the first-stage estimator by computing the functional derivative of the first-order condition with respect to the nuisance functions. This gives us a Riesz representer that, together with a stochastic equicontinuity property, enables us to express the risk-aversion parameter as an integral functional of the quantile of winning bids and its derivative. This particular representation allows us to invoke Theorem 4 of GG22 to establish a CLT. Our analysis otherwise follows the general procedure to obtain the limiting distribution of a semiparametric extremum estimator, for example, see Newey, (1994), Chen et al., (2003), and Ichimura and Lee, (2010).
The rest of the paper is organized as follows. Section 2 introduces the model, establishes identification of bidder valuations, and characterizes optimal reserve-price behavior under risk aversion. Section 3 develops the estimation methodology and the asymptotic properties of the AQR estimators for observed bids and bidder valuations. Section 4 presents the estimation of the risk-aversion parameter and derives its asymptotic properties. Section 5 reports a Monte Carlo study of the estimators. Section 6 provides an empirical application to real estate auctions in São Paulo. Section 7 concludes. Proofs of the main results are collected in the Appendix.
2 Model and identification
Consider an ascending auction of an indivisible object with $I\geq 2$ bidders. Bidders can raise prices continuously and without cost until only one bidder remains. The object is thus sold to the highest bidder for the price of his last bid, provided that it is at least as high as the reservation price. We assume an independent private values (IPV) environment, where each bidder knows only their own valuation and values are independently drawn across bidders.
Let the auctioned object have observable characteristics $X\in\mathbb{R}^{D}$. Bidder $i$’s valuation of the object is $V_{i}$, taking value in $\mathcal{V}=\left[\underline{v},\overline{v}\right]$. For notational simplicity, we assume $\mathcal{V}$ to be independent of $X$. We denote the conditional CDF of $V_{i}$ by $F\left(\cdot|X\right)$. To facilitate readers, as it may be instructive to compare our assumptions and results with GG22, we try to adopt their notations and termiologies when possible in this and the next section.
2.1 Bidders’ behavior
The winning bid, denoted by $B=V_{i}^{I-1:I}\in\mathcal{V}$, is the $\left(I-1\right)$-th order statistic among the $I$ i.i.d. private values $\left\{V_{i}\right\}_{i=1}^{I}$. This is the same equilibrium play for ascending auctions as used in Aradillas-López et al., (2013) and Gimenes, (2017). We denote the conditional CDF of $B$ by $G\left(\cdot|X\right)$.
M1(i) imposes a minimal regularity condition on $F\left(\cdot|X\right)$. M1(ii) is a structural assumption on the bidder’s bidding behavior, following Athey and Haile, (2002), described using the relation between the CDF of the $\left(I-1\right)$-th order statistic and the underlying CDF that the sample is drawn from.
Assumption M1.
(i) $F\left(\cdot|X\right)$ is differentiable and strictly increasing almost surely on $\mathcal{V}$;
(ii) $G\left(t|X\right)=\phi\left(F\left(t|X\right)\right)$ a.s. for $t\in\mathcal{V}$, where $\phi\left(a\right)=Ia^{I-1}-\left(I-1\right)a^{I}$ for $a\in\left[0,1\right]$.
Our paper takes a quantile approach to model the bidder’s private value distribution. Let the $\alpha$-quantile of the valuation distribution be denoted by $V\left(\alpha|X\right)=F^{-1}\left(\alpha|X\right)$ for $\alpha\in\left[0,1\right]$. As done in Gimenes, (2017) and GG22, we take $V_{i}=V\left(A_{i}|X\right)$ where $A_{i}\sim Uni\left[0,1\right]$ can be viewed as the private rank of the bidder, which is independent of $X$ and other bidders’ ranks.
We denote the $\alpha$-quantile of $B$ by $B\left(\alpha|X\right)=G^{-1}\left(\alpha|X\right)$ for $\alpha\in\left[0,1\right]$. $B\left(\cdot|X\right)$ exists, because $G\left(\cdot|X\right)$ is differentiable and strictly increasing, since $\phi:\left[0,1\right]\rightarrow\left[0,1\right]$ is continuously differentiable and is strictly increasing on $\left[0,1\right]$.
The quantile functions of the winning bid and the private valuation are linked through a one-to-one relationship. Proposition 1, which maps the quantile function of the private valuation from the quantile function of the winning bid, is the central identification argument on the bidder’s side (Gimenes, (2017)).
Proposition 1. Suppose Assumption M1 holds, then
$V(\alpha|X)=B\left(\phi(\alpha)|X\right)\text{\ a.s.\ for all }\alpha\in\left[0,1\right].$ (1)
2.2 Seller’s behavior
The seller can influence her expected revenue in an auction by setting reserve price. If all bids are below the reserve price, the seller keeps the object. If all but one bids are below the reserve price, the object is sold with the winner paying the reserve. If two or more bids are above the reserve price, the object is sold with the winner paying the second highest private value among all bidders.
The seller, whose value of the sale object is $W$, sets a reserve price, $R$, optimally to maximize her expected utility. Specifically, if the seller has utility function $U\left(\cdot\right)$, the expected utility from setting reserve price to be $r\in\mathcal{V}$ is:
$\widetilde{\Pi}(r,X,W)=U\left(W\right)F\left(r|X\right)^{I}+U\left(r\right)IF\left(r|X\right)^{I-1}\left(1-F\left(r|X\right)\right)+\int_{r}^{\overline{v}}U\left(t\right)dG\left(t|X\right).$ (2)
We assume $R=\arg\max_{{}_{r\in\mathcal{V}}}\widetilde{\Pi}\left(r,X,W\right)$, whose existence and uniqueness are guaranteed under the conditions of Assumption M2 given below.
Complementary to writing the bidders’ bids in terms of private ranks, the expected revenue can be equivalently expressed as a function of the screening level instead of the reserve price. For a screening level $\alpha$ that takes value in $\left[0,1\right]$, let us define
$\Pi(\alpha,X,W)=U\left(W\right)\alpha^{I}+U\left(V\left(\alpha|X\right)\right)I\alpha^{I-1}\left(1-\alpha\right)+I\left(I-1\right)\int_{\alpha}^{1}U\left(V\left(t|X\right)\right)t^{I-2}\left(1-t\right)dt,\text{ }$ (3)
so that $\Pi(\alpha,X,W)=\widetilde{\Pi}\left(r,X,W\right)$ when $r=V\left(\alpha|X\right)$. The optimal screening level is defined as $\alpha_{R}=\arg\max_{{}_{\alpha\in\left[0,1\right]}}\Pi\left(\alpha,X,W\right)$.
Assumption M2.
(i) $V_{i}$ has a conditional PDF, denoted by $f\left(\cdot|X\right)$, that is bounded away from zero and infinity a.s. on $\mathcal{V}$;
(ii) $J(v|X)=v-\frac{1-F(v|X)}{f\left(v|X\right)}$, defined for $v\in\mathcal{V}$, is strictly increasing a.s. on $\mathcal{V}$;
(iii) $W$ takes value in $\mathcal{W}\subseteq\mathcal{V}$;
(iv) $U\left(\cdot\right)$ is a twice continuously differentiable function with $U^{\left(1\right)}\left(\cdot\right)>0$ and $U^{\left(2\right)}\left(\cdot\right)\leq 0$.
Assumption M2 consists of standard conditions in the auction literature when studying the seller’s behavior. M2(i) is a regularity condition where the bounding from below ensures $J\left(\cdot\right)$ in M2(ii) is well defined. In his seminal paper, Myerson, (1981) calls $J\left(\cdot\right)$ the virtual valuation function, as it represents the marginal revenue contribution of a bidder with valuation $v$. He imposes monotonicity of $J\left(\cdot\right)$, which holds when $V_{i}$ has an increasing hazard rate, to prove incentive compatibility in the design of optimal auctions. In M2(iii), the lower bound on $W$ rules out point mass of the seller setting the reserve at $\underline{v}$, and the seller would be better off not selling the object if $W>\overline{v}$. M2(iv) assumes the utility function is smooth and allows the seller to be risk-averse as well as risk-neutral. The monotonicity and concavity of $U\left(\cdot\right)$ in M2(iv) are standard assumptions in the risk aversion literature where the differentiability conditions are imposed to facilitate analytical tractability – for example, it ensures Arrow-Pratt measure of risk aversion to be defined.
For two utility functions that are strictly increasing and weakly concave, $U_{1}(\cdot)$ and $U_{2}(\cdot)$, we say that $U_{2}(\cdot)$ represents a strictly more risk-averse preference in the Arrow-Pratt sense if there exists a real-valued function $\zeta(\cdot)$ that is twice continuously differentiable with $\zeta^{\left(1\right)}\left(\cdot\right)>0$ and $\zeta^{\left(2\right)}\left(\cdot\right)<0$ such that $U_{2}(\cdot)=\zeta(U_{1}(\cdot))$. [^3] It is not necessary to define Arrow-Pratt risk aversion with differentiable $\zeta(\cdot)$, and strict monotonicity and concavity will suffice. However, similarly to how we consider a smooth utility function in M2(iv), differentiability is used to facilitate analytical tractability. For differentiable utility functions, this formulation is equivalent to saying that their Arrow-Pratt risk aversions satisfy $\frac{U_{2}^{\left(2\right)}\left(v\right)}{U_{2}^{\left(1\right)}\left(v\right)}>-\frac{U_{1}^{\left(2\right)}\left(v\right)}{U_{1}^{\left(1\right)}\left(v\right)}$ for all $v$. The degree of risk aversion can take on a more compact form when utility functions belong to some parametric families. For example, when constant ARA (CARA) or RRA (CRRA) functions are used, ranking of Arrow-Pratt risk aversion between preferences is simply determined by the risk aversion parameters. For more background materials on Arrow-Pratt risk aversion, we refer the reader to Chapter 6.D in Mas-Colell et al., (1995).
Under Assumption M2, Proposition 2(i) below says that the optimal reserve price is characterized by the first-order condition obtained from differentiating (2), and Proposition 2(ii) says that the optimal reserve price decreases with the seller’s risk aversion in the Arrow-Pratt sense.
Proposition 2. Suppose Assumption M2 holds, then:
(i) For any $\left(X,W\right)$, the optimal reserve price exists and is uniquely determined by $r^{\ast}\in\mathcal{V}$ that satisfies
$0=U\left(W\right)+U^{\left(1\right)}\left(r^{\ast}\right)\frac{1-F(r^{\ast}|X)}{f\left(r^{\ast}|X\right)}-U\left(r^{\ast}\right);$ (4)
(ii) The optimal reserve price decreases with the seller’s risk aversion in the Arrow-Pratt sense.
The right hand side of equation (4) is the partial derivative of (2) with respect to $r$ evaluated at $r^{\ast}$. By putting $R$ in place of $r^{\ast}$ in (4) and use the identity that $R=V\left(\alpha_{R}|X\right)$, Proposition 2 confirms that $R$ is the unique maximizer of $\widetilde{\Pi}\left(\cdot,X,W\right)$ , and it satisfies
$0=U\left(W\right)+U^{\left(1\right)}\left(R\right)V^{\left(1\right)}\left(\alpha_{R}|X\right)\left(1-\alpha_{R}\right)-U\left(R\right).$ (5)
Later on, we will assume the shape of $U\left(\cdot\right)$ is known up to the risk aversion parameter. For example, in our empirical application, we use the CRRA utility function:
$U_{\theta}\left(v\right)=\left\{\begin{array}[]{c}\frac{v^{1-\theta}-1}{1-\theta}\\ \ln\left(v\right)\end{array}\begin{array}[]{c}\theta\neq 1\\ \theta=1\end{array}\right$., (6)
defined for $v>0$ and $\theta\in\mathbb{R}$, where higher $\theta$ represents a higher degree of risk aversion. Under the CRRA specification, M2(iv) holds for $\theta\geq 0$. We show in the proof of Lemma 4 how Proposition 2 can be used to identify an Arrow-Pratt risk aversion parameter.
3 Augmented quantile regression
Given data on winning bids and auction characteristics, this section proposes estimators for the quantile function of private values and related functions. We give the pointwise asymptotic properties of the quantile function estimator and the convergence rates for the estimators of the derivatives and inverse of the quantile function. The convergence rates are needed for deriving the large sample properties of the risk-aversion estimator in Section 4.
Our estimation strategy proceeds in two steps. We first estimate the quantile function of the winning bids. Then, we apply the identity linking bid and value quantiles given in Proposition 1. The first step follows what GG22 calls the augmented quantile regression (AQR) approach. Estimators for other functions of interest can be obtained subsequently. Note that seller’s specific variables, namely reserve price and outside value for the sale object, are not used for quantile estimation.
3.1 Assumptions for AQR estimation
We begin with some assumptions.
Assumption Q.
(i) The auction variables $\left\{\left(B_{l},X_{l}\right)\right\}_{l=1}^{L}$ are a random sample. For some $F\left(\cdot|X\right)$ and $G\left(\cdot|X\right)$ that satisfy Assumptions M1 and M2(i), $B_{l}$ takes value in $V$ and have conditional CDF $G\left(\cdot|X\right)$. $X_{l}$ takes value in $X\subseteq R^{D}$ such that $X$ is compact, and the eigenvalues of $E\left[X_{l}X_{l}^{\top}\right]$ are bounded away from zero and infinity.
(ii) Let $V\left(\alpha|X\right)=F^{-1}\left(\alpha|X\right)$ and $V\left(\alpha|X\right)=X_{1}^{\top}\gamma\left(\alpha\right)=\gamma_{0}\left(\alpha\right)+X^{\top}\gamma_{1}\left(\alpha\right)$ where $\gamma\left(\cdot\right)=\left[\gamma_{0}\left(\cdot\right),\gamma_{1}^{\top}\left(\cdot\right)\right]^{\top}$ is $\left(s+1\right)$-times continuously differentiable over $\left[0,1\right]$ for some $s\geq 1$ and $X_{1}=\left[1,X^{\top}\right]^{\top}$.
(iii) The kernel function $K\left(\cdot\right)$ is symmetric, continuously differentiable, and non-negative function on its support, $\left(-1,1\right)$. The bandwidth $h$ is positive and satisfy $h=o\left(1\right)$ and $\log^{2}L=o\left(Lh\right)$ as $L\rightarrow\infty$.
Q(i) imposes standard regularity conditions and correct model specification. In Q(ii), we assume that the quantile function of private values is linear in covariates and satisfies standard smoothness conditions. Q(iii) specifies the class of kernel functions and the conditions imposed on the bandwidth.
It is instructive to compare our assumptions with those found in Section 5.1 of GG22. First, we simplify their setting slightly by considering repeated auctions with a fixed number of bidders, while they allow the number of bidders to vary exogenously. It is straightforward for us to include this feature with more notation. Our Q(i) and Q(ii) are analogous to their Assumption A and Assumption S respectively. Our Q(iii) is the same with their Assumption H other than they require $\log^{2}L=o\left(Lh^{2}\right)$ as $L\rightarrow\infty$. GG22 imposes a more stringent requirement on the bandwidth than us, because we are studying different auction models. Specifically, we are estimating the quantile function of private value using winning bids from ascending auctions, whereas GG22 uses individual bids from first-price auctions. This matters, as the quantile function of the bidder’s private value depends on both the quantile function of the optimal bid and its derivative in a first price auction [^4] If $B\left(\cdot\right)$ and $B^{\left(1\right)}\left(\cdot\right)$ respectively were to denote the quantile function of the optimal first price bid. Then it can be shown that: $B\left(\alpha|X\right)$ $=$ $\frac{I-1}{\alpha^{I-1}}\int_{0}^{\alpha}a^{I-2}V\left(a|X\right)da\text{ with }\lim_{\alpha\rightarrow 0}B\left(\alpha|X\right)=V\left(0|X\right),$ $V\left(\alpha|X\right)$ $=$ $B\left(\alpha|X\right)+\frac{\alpha B^{\left(1\right)}\left(\alpha|X\right)}{I-1},$ see equations (2.4) and (2.5) in GG22., which contrasts with equation (1) where the quantile functions of the bidder’s private value and winning bids have the same degree of smoothness. Thus, GG22 cannot have the bandwidth decay too rapidly, since the variance of the derivative of quantile estimator is inversely proportional to the bandwidth while the bandwidth only appears in the higher order terms for the variance of the level quantile estimator. We will impose the same bandwidth condition as their Assumption H when we provide the convergence rates of the derivative of the quantile function.
3.2 Estimator of quantile function
The winning bid’s quantile function shares the linear quantile specification as the private value’s quantile function under Assumption Q. This follows from combining Q(ii) with the identity in (1), which gives:
$B\left(\alpha|X\right)=X_{1}^{\top}\beta\left(\alpha\right)\text{ \ with \ }\gamma\left(\alpha\right)=\beta\left(\phi(\alpha)\right)\text{ a.s. for }\alpha\in\left[0,1\right],$ (7)
recalling that $X_{1}=\left[1,X^{\top}\right]^{\top}$ so $\beta\left(\cdot\right)=\left[\beta_{0}\left(\cdot\right),\beta_{1}^{\top}\left(\cdot\right)\right]^{\top}$. Since $\gamma\left(\cdot\right)$ is a composite function of $\beta\left(\cdot\right)$ and $\phi\left(\cdot\right)$, estimating $V\left(\cdot\right)$ amounts to estimating $\beta\left(\cdot\right)$.
To motivate the AQR estimator for estimating $\beta\left(\cdot\right)$, first recall that
$B\left(\alpha|X\right)=\arg\min_{q}E\left[\rho_{\alpha}\left(B_{l}-q\right)|X\right]\text{ for }\alpha\in\left[0,1\right],$
where $\rho_{\alpha}\left(t\right)=t\left(\alpha-\mathbf{1}\left[t<0\right]\right)$ is the check function. The minimizer of the sample counterpart to the expectation above is the classic quantile regression estimator of Koenker and Bassett, (1978). This estimator is known to not perform well at extreme quantiles ($\alpha$ close to 0 or $1$). Moreover, it is piecewise linear and different estimators for quantile derivatives are required that complicates analysis of statistics that involve quantile level and its derivatives. Instead, let us consider $B\left(\cdot|X\right)$ over $\left[\alpha-h,\alpha+h\right]\cap\left[0,1\right]$, $\left\{B\left(\tau|X\right),\tau\in\left[\alpha-h,\alpha+h\right]\cap\left[0,1\right]\right\}$, which minimizes
$\int_{0}^{1}E\left[\rho_{a}\left(B_{l}-q\left(a,X\right)\right)|X\right]\frac{1}{h}K\left(\frac{a-\alpha}{h}\right)da,$
over any functions $q\left(\cdot,X\right)$ since $K\left(\cdot\right)$ is non-negative. In the same spirit as a local polynomial estimator (e.g., Fan and Gijbels, (1996)), the AQR approach, motivated by the Taylor’s expansion, estimates the quantile coefficients and their derivatives simultaneously. Specifically, with $B\left(\alpha+th|x\right)=\sum\limits_{j=0}^{s}x_{1}^{\top}\beta^{\left(j\right)}\left(\alpha\right)\frac{\left(th\right)^{j}}{j!}+O\left(h^{s+1}\right)$ in mind, consider the following objective function,
$\widehat{\mathcal{R}}\left(b;\alpha\right)$ $=$ $\frac{1}{L}\sum\limits_{l=1}^{L}\int_{0}^{1}\rho_{a}\left(B_{l}-P\left(X_{l},a-\alpha\right)^{\top}b\right)\frac{1}{h}K\left(\frac{a-\alpha}{h}\right)da$
$=$ $\frac{1}{L}\sum\limits_{l=1}^{L}\int_{-\frac{\alpha}{h}}^{\frac{1-\alpha}{h}}\rho_{\alpha+th}\left(B_{l}-P\left(X_{l},th\right)^{\top}b\right)K\left(t\right)dt,$
where $P\left(x,t\right)=\pi\left(t\right)\otimes x_{1}$ with $\pi\left(t\right)=\left[1,t,\ldots,\frac{t^{s}}{s!}\right]^{\top}$ and $x_{1}=\left[1,x^{\top}\right]^{\top}$. [^5] It is worth pointing out a subtle difference between our objective function relative to GG22 here. Despite of us both assuming $\left(s+1\right)$-times continuous differentiability of $V\left(\cdot\right)$, we make a polynomial approximation up to the $s$-th power term while GG22 goes up to the $\left(s+1\right)$-th power term. This is because the quantile function of the optimal first price auction bid has one more derivative than the quantile function of the private value – see equation (2.4) in GG22, which is given in the previous footnote.
Let $b\left(\alpha\right)=\left[\beta\left(\alpha\right)^{\top},\ldots,\beta^{\left(s\right)}\left(\alpha\right)^{\top}\right]^{\top}\in\mathbb{R}^{\left(s+1\right)\left(D+1\right)}$, so that $P\left(x,th\right)^{\top}b\left(\alpha\right)=\sum\limits_{j=0}^{s}x_{1}^{\top}\beta^{\left(j\right)}\left(\alpha\right)\frac{\left(th\right)^{j}}{j!}$. Our estimator of $b\left(\alpha\right)$ is $\widehat{b}\left(\alpha\right)=\arg\min_{b}\widehat{\mathcal{R}}\left(b;\alpha\right)$. We estimate $\beta\left(\alpha\right)$ by $\mathsf{S}_{0}\widehat{b}\left(\alpha\right)$, where $\mathsf{S}_{0}=S_{0}\otimes\mathrm{I}_{D+1}$ with $S_{0}=\left[1,\ldots,0\right]\in\mathbb{R}^{s+1}$, so that
$\widehat{B}\left(\alpha|x\right)=x_{1}^{\top}\mathsf{S}_{0}\widehat{b}\left(\alpha\right),$
is the estimator of $B\left(\alpha|x\right)$. We then estimate $V\left(\alpha|x\right)$ by using equation (1) in Proposition 1:
$\widehat{V}\left(\alpha|x\right)=\widehat{B}\left(\phi\left(\alpha\right)|x\right)\text{.}$ (8)
Before giving the statistical properties of $\widehat{V}\left(\alpha|x\right)$, we provide pointwise properties on $\widehat{B}\left(\alpha|x\right)$. This may be of independent interest for the AQR estimator, noting that these are not given in GG22.
Lemma 1. Suppose Assumption Q holds, there exists $\left(J_{B}\left(\alpha,x\right),J_{S}\left(\alpha,x\right),J_{R}\left(\alpha,x\right)\right)$ such that for all $\alpha\in\left(0,1\right)$ and $x\in\mathcal{X}$:
$\widehat{B}\left(\alpha|x\right)-B\left(\alpha|x\right)$ $=$ $J_{B}\left(\alpha,x\right)+J_{S}\left(\alpha,x\right)+J_{R}\left(\alpha,x\right),\text{ where}$
$J_{B}\left(\alpha,x\right)$ $=$ $h^{s+1}\mathsf{Bias}_{h}\left(\alpha,x\right)\mathsf{,}$
$E\left[J_{S}\left(\alpha,x\right)\right]$ $=$ $0\text{ and }\sqrt{L}J_{S}\left(\alpha,x\right)\overset{d}{\rightarrow}N\left(0,x_{1}^{\top}\Sigma\left(\alpha\right)x_{1}\right),$
$\sqrt{L}J_{R}\left(\alpha,x\right)$ $=$ $o_{p}\left(1\right),$
for $\underline{t}_{\alpha,h}=-\min\left(1,\frac{\alpha}{h}\right)$, $\overline{t}_{\alpha,h}=\max\left(1,\frac{1-\alpha}{h}\right)$,
$\mathsf{Bias}_{h}\left(\alpha,x\right)$ $=$ $B^{\left(s+1\right)}\left(\alpha|x\right)S_{0}\Omega_{h}\left(\alpha\right)^{-1}\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\frac{t^{s+1}\pi\left(t\right)}{\left(s+1\right)!}K\left(t\right)dt\text{ with}$
$\Omega_{h}\left(\alpha\right)$ $=$ $\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\pi\left(t\right)\pi\left(t\right)^{\top}K\left(t\right)dt,\text{ and }$
$\Sigma\left(\alpha\right)$ $=$ $\alpha\left(1-\alpha\right)\mathbf{P}_{0}\left(\alpha\right)^{-1}\mathbf{PP}_{0}\left(\alpha\right)^{-1}\text{ with}$
$\mathbf{P}_{0}\left(\alpha\right)$ $=$ $E\left[\frac{X_{l}X_{l}^{\top}}{B^{\left(1\right)}\left(\alpha|X_{l}\right)}\right]\text{ and }\mathbf{P}=E\left[X_{l}X_{l}^{\top}\right]\text{.}$
Moreover,
$\sup_{\left(\alpha,x\right)\in\left[0,1\right]\times\mathcal{X}}\left|\widehat{B}\left(\alpha|x\right)-B\left(\alpha|x\right)\right|=O_{p}\left(\sqrt{\frac{\log L}{L}}+h^{s+1}\right).$
The appendix gives expressions for the AQR estimator decomposition, $\left(J_{B}\left(\alpha,x\right),J_{S}\left(\alpha,x\right),J_{R}\left(\alpha,x\right)\right)$, in equations (18) to (20). These terms respectively represent the bias, leading stochastic term, and remainder term of $\widehat{B}\left(\alpha|x\right)-B\left(\alpha|x\right)$. Note that the limiting distribution of $\sqrt{L}J_{S}\left(\alpha,x\right)$ is the same as the standard quantile regression estimator’s without smoothing (for example, see Chapter 4 of Koenker, (2005)), so that the AQR estimator has the same first order asymptotic property as the Koenker and Bassett’s estimator when $Lh^{2\left(s+1\right)}=o\left(1\right)$.
[…]
Proposition 3. Suppose Assumption Q holds, for all $\alpha\in\left(0,1\right)$ and $x\in\mathcal{X}$:
$\widehat{V}\left(\alpha|x\right)-V\left(\alpha|x\right)=J_{B}\left(\phi\left(\alpha,x\right)\right)+J_{S}\left(\phi\left(\alpha,x\right)\right)+J_{R}\left(\phi\left(\alpha,x\right)\right),$
for the same functions $\left(J_{B}\left(\alpha,x\right),J_{S}\left(\alpha,x\right),J_{R}\left(\alpha,x\right)\right)$ as in Lemma 1. Moreover,
$\sup_{\left(\alpha,x\right)\in\left[0,1\right]\times\mathcal{X}}\left|\widehat{V}\left(\alpha|x\right)-V\left(\alpha|x\right)\right|=O_{p}\left(\sqrt{\frac{\log L}{L}}+h^{s+1}\right).$
3.3 Estimators of related functions
To prepare for the estimation of the risk parameter in the next section, we need to establish convergence rates for other functions related to the quantile. Let us re-write the first-order condition in (5) and replace $\alpha_{R}$ with $R=V^{-1}\left(\alpha_{R}|X\right)$, which gives:
$0=U\left(W\right)+U^{\left(1\right)}\left(R\right)V^{\left(1\right)}\left(V^{-1}\left(R|X\right)|X\right)\left(1-V^{-1}\left(R|X\right)\right)-U\left(R\right).$ (9)
Using the above equation for estimation requires estimators for $\left(V^{\left(1\right)}\left(\cdot\right),V^{-1}\left(\cdot\right)\right)$. Since the convergence rates for an estimator of $V^{\left(j\right)}\left(\cdot\right)$ may be generally useful for semiparametric estimation, we provide convergence rates for these objects as well. In what follows, we provide the relations between $\left(V^{\left(j\right)}\left(\cdot\right),V^{-1}\left(\cdot\right)\right)$ and $\left(B^{\left(j\right)}\left(\cdot\right),B^{-1}\left(\cdot\right)\right)$. Then, we define our estimators as transformations of the AQR estimators of $\left(B^{\left(j\right)}\left(\cdot\right),B^{-1}\left(\cdot\right)\right)$ and give their rates of convergence. As done for the quantile function previously, it will be instructive to first provide some statistical properties of $\left(\widehat{B}^{\left(j\right)}\left(\cdot\right),\widehat{B}^{-1}\left(\cdot\right)\right)$ as lemmas.
3.3.1 Derivatives of the quantile function
We can differentiate (1) repeatedly to obtain the relationship between the derivatives of the winning bid’s and private value’s quantile function. While the relations are visually compact for lower order derivatives, such as
$V^{\left(1\right)}(\alpha|X)$ $=$ $\phi^{\left(1\right)}(\alpha)B^{\left(1\right)}\left(\phi(\alpha)|X\right),$
$V^{\left(2\right)}(\alpha|X)$ $=$ $\left(\phi^{\left(1\right)}(\alpha)\right)^{2}B^{\left(2\right)}\left(\phi(\alpha)|X\right)+\phi^{\left(2\right)}(\alpha)B^{\left(1\right)}\left(\phi(\alpha)|X\right),$
which hold a.s. for all $\alpha\in\left[0,1\right]$, it gets cumbersome quickly for higher derivatives. Since we are only providing the uniform convergence rates rather than pointwise properties of $V^{\left(j\right)}(\cdot)$, it suffices to know that:
$V^{\left(j\right)}(\alpha|X)=\sum_{k=1}^{j}B^{\left(k\right)}\left(\phi(\alpha)|X\right)\mathcal{J}_{kj}\left(\phi^{\left(1\right)}(\alpha),\phi^{\left(2\right)}(\alpha),\ldots,\phi^{\left(j-k+1\right)}(\alpha)\right)\text{\ a.s. for all }\alpha\in\left[0,1\right],$ (10)
where $\mathcal{J}_{kj}$ is a known continuous function that is uniformly bounded over $\left[0,1\right]$ for all $k$ and jj. [^6] This can be obtained by applying the Faà di Bruno’s formula for computing chain rule to higher derivatives, for the $j$-th derivative, and $\mathcal{J}_{kj}\left(\cdot\right)$ is the exponential Bell polynomial.
Since the AQR approach readily gives estimator for $B^{\left(j\right)}(\alpha|x)$ for $j=1,\ldots,s$, we estimate $V^{\left(j\right)}(\alpha|x)$ by
$\widehat{V}^{\left(j\right)}(\alpha|x)$ $=$ $\sum_{k=1}^{j}\widehat{B}^{\left(k\right)}\left(\phi(\alpha)|x\right)\mathcal{J}_{kj}\left(\phi^{\left(1\right)}(\alpha),\phi^{\left(2\right)}(\alpha),\ldots,\phi^{\left(j-k+1\right)}(\alpha)\right),\text{\ where}$
$\widehat{B}^{\left(j\right)}\left(\alpha|x\right)$ $=$ $x_{1}^{\top}\mathsf{S}_{j}\widehat{b}\left(\alpha\right),\text{\ }$
with $\mathsf{S}_{j}=S_{j}\otimes\mathrm{I}_{D+1}$ and $S_{j}$ is a row vector of size $\left(s+1\right)$ consists of 0’s in every component other than $1$ in its $\left(j+1\right)$-th entry.
Lemma 2 gives the convergence rate for the AQR estimator of $\widehat{B}^{\left(j\right)}\left(\cdot\right)$.
Lemma 2. Suppose Assumption Q holds and $\lim_{L\rightarrow\infty}\frac{\log^{2}L}{Lh^{2}}=0$, then for $j=1,2,\ldots,s$:
$\sup_{\left(\alpha,x\right)\in\left[0,1\right]\times\mathcal{X}}\left|\widehat{B}^{\left(j\right)}\left(\alpha|x\right)-B^{\left(j\right)}\left(\alpha|x\right)\right|=O_{p}\left(\sqrt{\frac{\log L}{Lh^{2j-1}}}+h^{s+1-j}\right).$
Notice that Lemma 2 imposes the same bandwidth condition as GG22, which we alluded earlier. Indeed, GG22 has given the same convergence rate as the above when $j=1$, see equation (5.7) in their Theorem 2.777The order of their bias is written as $h^{s+1}$, which is a result of their bid’s quantile function having one more derivative than ours. The component of the convergence rate that corresponds to the stochastic term is $\sqrt{\frac{\log L}{Lh^{2j-1}}}$, which coincides with the usual rates of the $\left(j-1\right)$-th derivative of a kernel density estimator; this finding is reassuring given the identity between the density and derivative of the quantile. The worsening of the bias rate with higher derivatives also mirrors standard local polynomial estimators. Since $\widehat{V}^{\left(j\right)}\left(\cdot\right)$ is a smooth in $\left\{\widehat{B}^{\left(k\right)}\left(\cdot\right)\right\}_{k=1}^{j}$, its rate of convergence then follows that of $\widehat{B}^{\left(j\right)}\left(\cdot\right)$.
Proposition 4. Suppose Assumption Q holds and $\lim_{L\rightarrow\infty}\frac{\log^{2}L}{Lh^{2}}=0$, then for $j=1,2,\ldots,s$:
$\sup_{\left(\alpha,x\right)\in\left[0,1\right]\times\mathcal{X}}\left|\widehat{V}^{\left(j\right)}\left(\alpha|x\right)-V^{\left(j\right)}\left(\alpha|x\right)\right|=O_{p}\left(\sqrt{\frac{\log L}{Lh^{2j-1}}}+h^{s+1-j}\right).$
3.3.2 Inverse of the quantile function
The inverse of the quantile function is the CDF. The relation between the inverse of private value and winning bid quantiles are given by the inverse of composite functions formula applied to (1):
$V^{-1}(t|X)=\phi^{-1}(B^{-1}\left(t|X\right))\text{\ \ a.s. for all }t\in\mathcal{V}\text{.}$ (11)
We define our estimator for $V^{-1}\left(\cdot\right)$ as follows,
$\widehat{V}^{-1}(t|x)$ $=$ $\phi^{-1}(\widehat{B}^{-1}\left(t|x\right)),\text{ for all }\left(t,x\right)\in\mathcal{V}\times\mathcal{X}\text{ where,}$
$\widehat{B}^{-1}\left(t|x\right)$ $=$ $\inf\left\{a\in\left[0,1\right]:\widehat{B}\left(a|x\right)\geq t\right\}.$
The statistical properties of $\widehat{V}^{-1}(t|x)$ can be analyzed through two applications of the continuous mapping theorem.
First, $\widehat{B}^{-1}\left(t|x\right)$ can be studied as the inverse of $\widehat{B}\left(t|x\right)$, through the map $\Psi:\ell^{\infty}\left(\left[0,1\right]\right)\mapsto\ell^{\infty}\left(\mathcal{V}\right)$, such that $\Psi\left(\pi\right)\left(t\right)=\inf\left\{a\in\left[0,1\right]:\pi\left(a\right)\geq t\right\}$ for $t\in\mathcal{V}$ and $\pi\left(\cdot\right)\in\ell^{\infty}\left(\left[0,1\right]\right)$. Here we use $\ell^{\infty}\left(\mathcal{A}\right)$ to denote the space of bounded functions on $\mathcal{A}\subseteq\mathbb{R}$. It is well known that the Hadamard derivative of $\Psi$ exists and linearization methods apply (van der Vaart and Wellner, (2023), Lemma 3.10.21). Particularly, the leading term from linearizing $\widehat{B}^{-1}\left(t|x\right)-B^{-1}\left(t|x\right)$ is,
$\frac{\widehat{B}\left(B^{-1}\left(t|x\right)|x\right)-B\left(B^{-1}\left(t|x\right)|x\right)}{B^{\left(1\right)}\left(B^{-1}\left(t|x\right)|x\right)},$ (12)
when $B^{\left(1\right)}\left(B^{-1}\left(t|x\right)|x\right)>0$.
Lemma 3 gives, along side the uniform convergence rate, the pointwise properties of $\widehat{B}^{-1}\left(t|x\right)$ for $t\in int\left(\mathcal{V}\right)$, denoting in the interior of $\mathcal{V}$. The latter may be of independent interest, as $\widehat{B}^{-1}\left(t|x\right)$ is a flexible yet low-dimensional general estimator for the conditional CDF.
Lemma 3. Suppose Assumption Q holds, for all $t\in int\left(\mathcal{V}\right)$ and $x\in\mathcal{X}$:
$\widehat{B}^{-1}\left(t|x\right)-B^{-1}\left(t|x\right)$ $=$ $H_{B}\left(t,x\right)+H_{S}\left(t,x\right)+H_{R}\left(t,x\right),\text{ where}$
$H_{B}\left(t,x\right)$ $=$ $-\frac{h^{s+1}}{B^{\left(1\right)}\left(B^{-1}\left(t|x\right)|x\right)}\mathsf{Bias}_{h}\left(B^{-1}\left(t|x\right),x\right),$
$E\left[H_{S}\left(t,x\right)\right]$ $=$ $0\text{ and }\sqrt{L}H_{S}\left(t,x\right)\overset{d}{\rightarrow}N\left(0,\frac{x_{1}^{\top}\Sigma\left(B^{-1}\left(t|x\right)\right)x_{1}}{\left(B^{\left(1\right)}\left(B^{-1}\left(t|x\right)|x\right)\right)^{2}}\right),$
$\sqrt{L}H_{R}\left(t,x\right)$ $=$ $o_{p}\left(1\right),$
for the same $\mathsf{Bias}_{h}\left(\cdot\right)$ and $\Sigma\left(\cdot\right)$ as in Lemma 1. Moreover,
$\sup_{\left(t,x\right)\in\mathcal{V}\times\mathcal{X}}\left|\widehat{B}^{-1}\left(t|x\right)-B^{-1}\left(t|x\right)\right|=O_{p}\left(\sqrt{\frac{\log L}{L}}+h^{s+1}\right).$
$\phi\left(a\right)$ $=$ $Ia^{I-1}-\left(I-1\right)a^{I}$
$\phi^{\prime}\left(a\right)$ $=$ $I\left(I-1\right)a^{I-2}\left(1-I\left(I-1\right)a\right)$
Second, we apply $\phi^{-1}\left(\cdot\right)$ to $\widehat{B}^{-1}\left(\cdot\right)$. There is a potential complication when deriving uniform convergence in this step as $\phi^{-1}\left(\cdot\right)$ is continuously differentiable on $\left(0,1\right)$ but not at the boundaries. This can be seen from inspecting the derivative of $\phi^{-1}\left(\cdot\right)$, which is $1/\phi^{\left(1\right)}\left(\phi^{-1}\left(\cdot\right)\right)$, as we have $\phi^{\left(1\right)}\left(0\right)=0$ when $I>2$ and $\phi^{\left(1\right)}\left(1\right)$ is 0 for all $I$. It should be noted too that the bias and variance of $\widehat{B}^{-1}\left(\cdot\right)$ go to zero as $t\rightarrow\underline{v}$, and also $t\rightarrow\overline{v}$ for the variance. These faster convergence rates may mitigate potential irregularity issues for the de-meaned component of $\widehat{V}^{-1}\left(\cdot\right)$ as well as the lower boundary bias. Nevertheless, we do not need these aspects to derive the large sample properties of our risk aversion parameter in the next section, and a comprehensive study on uniform properties of such transformed AQR estimator is beyond the scope of this paper.
The next proposition gives the uniform convergence rate for $\widehat{V}^{-1}\left(\cdot\right)$ on an inner interval of $\mathcal{V}$, denoted by $\mathcal{V}_{\delta}$ and defined as $\left[\underline{v}+\delta,\overline{v}-\delta\right]$ for $\delta\in\left(0,\left(\overline{v}-\underline{v}\right)/2\right)$.
Proposition 5. Suppose Assumption Q holds, then
$\sup_{\left(t,x\right)\in\mathcal{V}_{\delta}\times\mathcal{X}}\left|\widehat{V}^{-1}\left(t|x\right)-V^{-1}\left(t|x\right)\right|=O_{p}\left(\sqrt{\frac{\log L}{L}}+h^{s+1}\right).$
4 Risk aversion estimator
We now assume the utility function takes a parametric form: $\theta\mapsto U_{\theta}\left(\cdot\right)$, for some $\theta\in\Theta\subset\mathbb{R}$ that represents an Arrow-Pratt measure of risk aversion. We can then construct an objective function for estimating the risk parameter from the first-order condition in (9). To do this, let us use $\mathcal{D}_{1}\left(\mathcal{A_{\delta}}\times\mathcal{X}\right)$ and $\mathcal{D}_{2}\left(\mathcal{V_{\delta}}\times\mathcal{X}\right)$ to denote classes of functions whose images are $\mathcal{V}_{\delta}$ and $\mathcal{A}_{\delta}$ respectively. We denote candidates for the derivative and inverse of the conditional quantile function of $V_{l}$ given $X_{l}$ by $\psi_{1}\left(\cdot\right)\in\mathcal{D}_{1}\left(\mathcal{A_{\delta}}\times\mathcal{X}\right)$ and $\psi_{2}\left(\cdot\right)\in\mathcal{D}_{2}\left(\mathcal{V_{\delta}}\times\mathcal{X}\right)$ respectively. Here, we use $\mathcal{V}_{\delta}=\left[\underline{v}+\delta,\overline{v}-\delta\right]$ and $\mathcal{A}_{\delta}=\left[\delta,1-\delta\right]$ for small $\delta$. We restrict the support of the quantile and valuation for the reason discussed at the end of Section 3. Note that $\delta$ in $\mathcal{V}_{\delta}$ and $\mathcal{A}_{\delta}$ can generally be different. Moreover, the supports of (bidder’s and seller’s) valuation (and the reserve price) can depend on $X$. Incorperating these is conceptually straightforward. We forego the more general notations for simplicity of presentation.
Consider the following real value function $q\left(z,\theta,\psi\right)$ defined as follows:
$q\left(z,\theta,\psi\right)=U_{\theta}\left(w\right)+U_{\theta}^{\left(1\right)}\left(r\right)\psi_{1}\left(\psi_{2}\left(r,x\right),x\right)\left(1-\psi_{2}\left(r,x\right)\right)-U_{\theta}\left(r\right),$ (13)
where $z=\left(w,r,x\right)\in\mathcal{Z}=\mathcal{W}\times\mathcal{V}_{\delta}\times\mathcal{X}$, $\theta\in\Theta$, and $\psi\left(\cdot\right)=\left(\psi_{1}\left(\cdot\right),\psi_{2}\left(\cdot\right)\right)\in\mathcal{D}_{1}\left(\mathcal{A_{\delta}}\times\mathcal{X}\right)\times\mathcal{D}_{2}\left(\mathcal{V_{\delta}}\times\mathcal{X}\right)$. Henceforth, we compress the arguments of functions that are parameters, i.e., $\psi\left(\cdot\right)$ to $\psi$, for notational brevity.
Our structural assumption will, by design, require that the first order condition in (9) coincides with $q\left(z,\theta_{0},\psi_{0}\right)$ for some $\left(\theta_{0},\psi_{0}\right)$ for all $z$ that is consistent with the auction model describe in Section 2. This is suggestive for a least squares or minimum distance type objective function for estimating $\theta_{0}$.
Let $P_{Z}$ denote a probability distribution of $Z=\left(W,R,X\right)$ and suppose $\left\{Z_{l}\right\}_{l=1}^{L}$ is a random sample drawn from it. We define,
$Q_{L}\left(\theta,\psi\right)=\frac{1}{L}\sum\limits_{l=1}^{L}q\left(Z_{l},\theta,\psi\right)^{2}\text{ \ and \ }Q\left(\theta,\psi\right)=\int q\left(Z,\theta,\psi\right)^{2}dP_{Z}\text{.}$
Given our usage of empirical process methods to proving the asymptotic results, we write $Q$ as an integral, making clear that only $Z$ is being integrated out, so that $Q\left(\theta,\psi\right)$ is a random variable if either $\theta\$or $\psi$ is random. We add that, more precisely, $P_{Z}$ can be understood as a conditional distribution for $R\in\mathcal{V}_{\delta}$ for the purpose of the proofs, although there is no data trucation in practice. Related to the latter point, we emphasize that, we are not at risk of identification loss with our minimum distance approach by working on $\left(\mathcal{A}_{\delta},\mathcal{V}_{\delta}\right)$ instead of $\left(\left[0,1\right],\mathcal{V}\right)$, which should be contrasted with choosing moments in a conditional moment model (e.g., see Domínguez and Lobato, (2004)), as a single value of $z\in\mathcal{Z}$ identifies $\theta_{0}$ via (13).
Our estimation problem here is a semiparametric one, as we are interested in the finite dimensional parameter $\theta_{0}$ in the presence of infinite dimensional nuisance functions. We denote the estimator for the latter by $\widehat{\psi}$, which consists of $\widehat{\psi}_{1}=\widehat{V}^{\left(1\right)}$ and $\widehat{\psi}_{2}=\widehat{V}^{-1}$, respectively defined as in 3.3.1 and 3.3.2 using a kernel function that satisfies Assumption Q(iii). We then define our estimator for $\theta_{0}$ to be any $\widehat{\theta}$ that satisfies:
$\left|Q_{L}\left(\widehat{\theta},\widehat{\psi}\right)\right|\leq\inf_{\theta\in\Theta}\left|Q_{L}\left(\theta,\widehat{\psi}\right)\right|+o_{p}\left(1/\sqrt{L}\right).$ (14)
We impose the following conditions.
Assumption S1.
(i) $q\left(Z,\theta_{0},\psi_{0}\right)=0$ $P_{Z}$-a.s. for some $\theta_{0}\in\Theta$, $\psi_{10}\in\mathcal{D}_{1}\left(\mathcal{A_{\delta}}\times\mathcal{X}\right)$, and $\psi_{20}\in\mathcal{D}_{2}\left(\mathcal{V_{\delta}}\times\mathcal{X}\right)$;
(ii) The auction variables $\left\{\left(B_{l},Z_{l}\right)\right\}_{l=1}^{L}$ are a random sample such that $Z_{l}\sim P_{Z}$, and the distribution of $\left(B_{l},Z_{l}\right)$ satisfies conditions in Assumptions M1, M2(i), M2(ii), M2(iii), and Q(i);
(iii) $\Theta$ is compact, $U_{\theta}$ is twice continuously differentiable with $U_{\theta}^{\left(1\right)}>0$ and $U_{\theta}^{\left(2\right)}\leq 0$ on $\mathcal{V}_{\delta}$ for all $\theta\in\Theta$ such that: (a) for any $\theta^{\prime}>\theta$, there exists $\zeta$ such that $\zeta^{\left(1\right)}>0$, $\zeta^{\left(2\right)}<0$ and $U_{\theta^{\prime}}=\zeta(U_{\theta})$; and (b) $\sup_{\theta\in\Theta}E\left[\left|U_{\theta}^{\left(j\right)}\left(R\right)\right|\right]<\infty$ for $j=0,1,2$;
(iv) $\mathcal{D}_{1}\left(\mathcal{A_{\delta}}\times\mathcal{X}\right)=\{$ $\frac{\partial}{\partial\alpha}\nu\left(\phi\left(\alpha\right),x\right)$ for $\nu\in\mathcal{D}_{0}\left(\mathcal{A_{\delta}}\times\mathcal{X}\right)$ }\} and $\mathcal{D}_{2}\left(\mathcal{V_{\delta}}\times\mathcal{X}\right)=\{$ $\left(t,x\right)\mapsto\nu_{x}^{-1}\left(t\right)$ for $\left(t,x\right)\in\mathcal{V}_{\delta}\times\mathcal{X}$ where $\nu_{x}\left(\alpha\right)=\nu\left(\phi\left(\alpha\right),x\right)$ for $\nu\in\mathcal{D}_{0}\left(\mathcal{A_{\delta}}\times\mathcal{X}\right)$ }\} where $\mathcal{D}_{0}\left(\mathcal{A_{\delta}}\times\mathcal{X}\right)=\{$ $x_{1}^{\top}\mu\left(\alpha\right)$ for $x\in\mathcal{X}$ and $\mu:\mathcal{A}_{\delta}\rightarrow\mathbb{R}$ is $\left(s+1\right)$-times continuously differentiable for some $s\geq 1$ such that: (a) $x_{1}^{\top}\mu\left(\alpha\right)$ takes value in $\mathcal{V}_{\delta}$ and is strictly increasing in $\alpha$; (b) $x_{1}^{\top}\mu^{\left(1\right)}\left(\alpha\right)$ is bounded away from zero and from infinity uniformly; and (c) $x_{1}^{\top}\mu^{\left(2\right)}\left(\alpha\right)$ is bounded away from infinity uniformly }\}, and the $\phi\left(\alpha\right)$-th quantile function of $B_{l}$ conditional on $X_{l}$ for $\phi\left(\alpha\right)\in\mathcal{A}_{\delta}$ lies in $\mathcal{D}_{0}\left(\mathcal{A}_{\delta}\times\mathcal{X}\right)$.
Assumption S1 consists of structural assumptions and regularity conditions on the variables in the model.
The condition $q\left(Z,\theta_{0},\psi_{0}\right)=0$ in S1(i) assumes correct model specification and, together with S1(ii), they imply the distribution of the data can be rationalized by the ascending auction model described in Section 2. Particularly, it implies that $\psi_{10}\left(\alpha,x\right)=V^{\left(1\right)}\left(\alpha|x\right)=\phi^{\left(1\right)}\left(\alpha\right)B^{\left(1\right)}\left(\phi\left(\alpha\right)|x\right)$ for $\left(\alpha,x\right)\in\mathcal{A}_{\delta}\times\mathcal{X}$ and $\psi_{20}\left(t,x\right)=V^{-1}\left(t|x\right)=\phi^{-1}\left(t\right)B^{-1}\left(\phi\left(t\right)|x\right)$ for $\left(t,x\right)\in\mathcal{V}_{\delta}\times\mathcal{X}$.
S1(iii) imposes a parametric assumption on the utility function that satisfies M2(iv). Part (a) gives an interpretation for $\theta$ to be an Arrow–Pratt coefficient where higher $\theta$ means higher degree of risk aversion in the Arrow–Pratt sense as described in Section 2.2. For example, the risk parameters in CARA and CRRA utility functions satisfy this condition (Ross, (1981)). Part (b) is a regularity condition requiring uniform squared integrability. In our application we use the CRRA utility function, which is defined in (6). It should be noted that such utility function is defined for all $\theta\in\mathbb{R}$, where $\theta>0$, $\theta=0,$ and $\theta<0$ represent risk-averse, risk-neutral, and risk-loving preferences respectively. Under the CRRA preference, S1(iii) requires $\theta\geq 0$. The sole purpose of this is to ensure we can apply Proposition 2, which we use to prove Lemma 4 that shows $\theta_{0}$ is identified as the minimizer of $Q$ under primitive conditions. Our estimation procedure does not restrict the parameter space to be non-negative. Importantly, if the implication of Lemma 4 is assumed to hold as a high-level identification condition, the asymptotic theory for our estimator applies for $\theta_{0}<0$ without any modification. [^8] Our simulation study does not suggest any identification issue when $\theta_{0}<0$, and the estimator in the risk-loving case behaves in the same was as the risk-neutral and risk-averse cases qualitatively. These additional simulation results are available upon request. We can therefore abstract away from the inference issues that arise when parameter is on the boundary such as those discussed in Andrews, (2001). Lastly, on parametric modelling, we can also embed observables in the risk measure, for example, by replacing $\theta$ with a linear index of the seller’s characteristics.
S1(iv) ensures $\mathcal{D}_{1}\left(\mathcal{A}_{\delta}\times\mathcal{X}\right)$ and $\mathcal{D}_{2}\left(\mathcal{V_{\delta}}\times\mathcal{X}\right)$ are the correct classes of functions containing candidates of derivative and inverse of quantile functions respectively. These classes of functions are derived from $\mathcal{D}_{0}\left(\mathcal{A_{\delta}}\times\mathcal{X}\right)$ can represent quantile functions that are linear in the regressors with additional regularity conditions: Part (a) ensures quantile inverses (i.e., CDFs) exist with image in $\mathcal{A}_{\delta}$; Part (b) requires the first derivatives of quantile functions are bounded away from zero and infinity, ensuring the corresponding PDFs satisfy Assumption M2(i); Part (c) imposes additional condition to ensure $\mathcal{Q}$ is a Glivenko-Cantelli class of functions under $P_{Z}$.
We note that the uniform boundedness conditions imposed on $\mathcal{D}_{0}\left(\mathcal{A}_{\delta}\times\mathcal{X}\right)$, and subsequently inherited by $\mathcal{D}_{1}\left(\mathcal{A}_{\delta}\times\mathcal{X}\right)$ and $\mathcal{D}_{2}\left(\mathcal{V_{\delta}}\times\mathcal{X}\right)$, are very mild in practice. This is because the true quantile function satisfies these conditions, and we have consistent estimators for the derivative and inverse of the quantile function. We therefore only need to consider $\mathcal{D}_{1}\left(\mathcal{A}_{\delta}\times\mathcal{X}\right)$ and $\mathcal{D}_{2}\left(\mathcal{V_{\delta}}\times\mathcal{X}\right)$ that contain functions in a neighborhood of $\left(\psi_{10},\psi_{20}\right)$. In a similar vein, the requirement for the image of functions in $\mathcal{D}_{0}\left(\mathcal{A}_{\delta}\times\mathcal{X}\right)$ and $\mathcal{D}_{2}\left(\mathcal{V}_{\delta}\times\mathcal{X}\right)$ to respectively be $\mathcal{V}_{\delta}$ and $\mathcal{A}_{\delta}$ is not restrictive for a fixed $\delta$, as our estimators for $\left(\psi_{10},\psi_{20}\right)$ converge to the true functions uniformly over any inner subset of their respective supports. Given the boundedness of functions involved, we shall use $\left\|\cdot\right\|_{\infty}$ to generically denote the sup-norm of functions over their supports.
Building on the result of Proposition 2, Lemma 4 says our population objective function has a well separated minimum at $\theta_{0}$ when $\psi=\psi_{0}$ under S1.
Lemma 4. Suppose Assumption S1 holds, then for all $\epsilon>0$, there exists $\delta>0$ such that $\inf_{\left|\theta-\theta_{0}\right|>\epsilon}Q\left(\theta,\psi_{0}\right)\geq Q\left(\theta_{0},\psi_{0}\right)+\delta$.
[…]
Lemma 5. Suppose Assumption S1 holds and the bandwidth satisfies $h=o\left(1\right)$ and $\log^{2}L=o\left(Lh^{2}\right)$, then $\sup_{\theta\in\Theta}\left|Q_{L}\left(\theta,\widehat{\psi}\right)-Q\left(\theta,\psi_{0}\right)\right|=o_{p}\left(1\right)$.
Theorem 1. Suppose Assumption S1 holds and the bandwidth satisfies $h=o\left(1\right)$ and $\log^{2}L=o\left(Lh^{2}\right)$, then $\widehat{\theta}=\theta_{0}+o_{p}\left(1\right)$.
[…]
Assumption S2.
(i) $U_{\theta}$ is three times continuously differentiable with $\sup_{\theta\in\Theta}E\left[\left|U_{\theta}^{\left(j\right)}\left(R\right)\right|^{2}\right]<\infty$ for $j=0,1,2,3$;
(ii) $\mathcal{D}_{0}\left(\mathcal{A}_{\delta}\times\mathcal{X}\right)$ is as described in S1(iv) other than $\mu$ is $\left(s+2\right)$-times continuously differentiable for some $s\geq 1$ and there exists an enveloping function for $\left(\alpha,x\right)\mapsto x_{1}^{\top}\mu^{\left(3\right)}\left(\alpha\right)$ that is $L_{2}\left(P_{Z_{0}}\right)$-integrable;
(iii) $E\left[\frac{\partial}{\partial\theta}q\left(Z_{l},\theta_{0},\psi_{0}\right)^{2}\right]$ is invertible.
The strengthened smoothness and moment conditions in S2(i) and S2(ii) ensure that $\mathcal{Q}$ and the related class of functions are $P_{Z}$-Donsker and allow us to bound various moments in the proof. S2(iii) is the invertible Hessian condition.
[…]
$\int c_{Y_{1}}\left(\widehat{\psi}_{1}-\psi_{10}\right)\circ\psi_{20}+c_{Y_{2}}\left(\widehat{\psi}_{2}-\psi_{20}\right)dP_{Z}+o_{p}\left(\frac{1}{\sqrt{L}}\right)\text{ where,}$
$\sqrt{L}\int c_{Y_{1}}\left(\widehat{\psi}_{1}-\psi_{10}\right)\circ\psi_{20}+c_{Y_{2}}\left(\widehat{\psi}_{2}-\psi_{20}\right)dP_{Z}\overset{d}{\rightarrow}N\left(0,\sigma_{0}^{2}\right)\text{ for some }\sigma_{0}^{2}\left.>\right.0\text{,}$
where $c_{Y_{1}}$ and $c_{Y_{2}}$ are functions of $Z$, and $\circ$ denotes the composition of functions. See (28) and (29) in Appendix C respectively for the explicit forms of $c_{Y_{1}}$ and $c_{Y_{2}}$.
[…]
It is worth noting that our bandwidth requirement of $\log^{2}L=o\left(Lh^{3}\right)$ coincides with the condition in Theorem 4 of GG22, which establishes asymptotic normality for a similar integral functional that contains the bid’s quantile function and its derivative. This is reassuring, as they estimate their functional using AQR estimators that have the same convergence rates as $\widehat{\psi}_{20}$ and $\widehat{\psi}_{10}$Ḟor example, when the bandwidth decays at the rate $\log L^{1/2}L^{-\varsigma}$ for some $\varsigma>0$, Lemma 6 requires $\frac{1}{4s}<\varsigma<\frac{1}{3}$ for it to hold.
Theorem 2. Suppose Assumptions S1 and S2 hold, the bandwidth satisfies $Lh^{4s}=o\left(1\right)$ and $\log^{2}L=o\left(Lh^{3}\right)$, and for the same $\sigma_{0}^{2}$ in Lemma 6,
$\sqrt{L}\left(\widehat{\theta}-\theta_{0}\right)\overset{d}{\rightarrow}N\left(0,\frac{\sigma_{0}^{2}}{\left(E\left[\frac{\partial}{\partial\theta}q\left(Z_{l},\theta_{0},\psi_{0}\right)^{2}\right]\right)^{2}}\right).$
In practice, we recommend that inference on $\theta_{0}$ be performed using the ordinary nonparametric bootstrap. Chen et al., (2003) provide high-level conditions under which the asymptotic distribution of a two-step semiparametric estimator can be consistently bootstrapped. We conjecture that our setup is amenable to such a result. Indeed, we apply this in our Monte Carlo study and find that bootstrap standard errors for $\widehat{\theta}$ perform very well. A formal proof of bootstrap validity when AQR estimators are used to estimate nuisance functions in a general two-step semiparametric procedure is, however, beyond the scope of this paper.
5 Simulation
In this section, we consider some finite sample properties of our AQR and semiparametric estimators proposed in the paper.
5.1 Simulation design
Taking inspirations from Gimenes, (2017) and GG22, the private-value quantile function is specified as follows:
$V(\alpha|X)$ $=$ $\gamma_{0}(\alpha)+\gamma_{1}(\alpha)X_{1}+\gamma_{2}(\alpha)X_{2}\text{, where}$
$\gamma_{0}(\alpha)$ $=$ $-\log(1-(1-1/e)\alpha),$
$\gamma_{1}(\alpha)$ $=$ $1,$
$\gamma_{2}(\alpha)$ $=$ $1-\exp(-5\alpha).$
The quantile functions $\gamma_{0}$, $\gamma_{1}$, $\gamma_{2}$ are all strictly increasing, while $\gamma_{0}$ is convex, $\gamma_{1}$ is linear, and $\gamma_{2}$ is concave. The covariates $X_{1}$ and $X_{2}$ are independent and uniformly distributed on $[0,1]$. For the outside option, we let $W(X)=V(\beta|X)$, where $\beta$ is an independent draw from a uniform distribution on $[0.05,0.5]$.
We use the CRRA utility function and consider $\theta_{0}=0,0.5,1$. Under this specification, $\theta_{0}>0$ means the seller is risk-averse and $\theta_{0}=0$ means the seller is risk-neutral (and $\theta_{0}<0$ means the seller is risk-loving).
We consider $L$ auctions with $3$ bidders, where $L=250,500,1000$. In the estimation, AQR quantile functions are computed over the estimation grid of $\alpha$ taking values 0.02,0.04,,0.980.02,0.04,...,0.98. Following GG22, we set the AQR polynomial order to be 2 and estimate it using the Epanechnikov kernel: $K(t)=0.75(1-t^{2})\mathbf{1}(t\in[-1,1])$. Three different bandwidths are used in the simulation: $h=sL^{-1/5},sL^{-1/6},sL^{-1/7}$, where $s$ is the sample standard deviation of the winning bids. The number of replications is 1000 in all experiments.
5.2 Simulation results
Figure 1 collects the simulation results for the estimation of the private-value quantile function and its derivative for different $L$. The black solid line is the true function, the red dashed line is the mean of our estimator, and the red dotted lines represent its $2.5$th and $97.5$th percentiles. In all these figures, we set $X_{1}=X_{2}=0.5$. Here we only report the estimation results of $V(\alpha|X)$ and $V^{(1)}(\alpha|X)$ using $h=sL^{-1/6}$. The results using $h=sL^{-1/5}$ and $h=sL^{-1/7}$ are similar.
We can see our estimator of $V$, on average, lies very close to the true, within the $95\%$ central quantile interval, and with decreasing variance as $L$ increases. The estimator of $V^{(1)}$ shares these properties but has higher estimation error, which is what we expect from the theory. To illustrate the differences more quantitatively, we calculate the integrated mean squared error (IMSE) for $\widehat{V}$ and $\widehat{V}^{(1)}$ when $X_{1}=X_{2}=0.5$. These can be found in the Table 1.
Figure 1: Simulation results for $\widehat{V}$ and $\widehat{V}^{(1)}$.
Table 1: IMSE for $\widehat{V}$ and $\widehat{V}^{(1)}$.
$h$ $L$ IMSE for $\widehat{V}$ IMSE for $\widehat{V}^{(1)}$
$sL^{-1/5}$ 250 $8.88\times 10^{-4}$ 0.0808
500 $4.19\times 10^{-4}$ 0.0475
1000 $2.13\times 10^{-4}$ 0.0315
$sL^{-1/6}$ 250 $8.85\times 10^{-4}$ 0.0733
500 $4.19\times 10^{-4}$ 0.0422
1000 $2.11\times 10^{-4}$ 0.0267
$sL^{-1/7}$ 250 $8.84\times 10^{-4}$ 0.0694
500 $4.21\times 10^{-4}$ 0.0391
1000 $2.13\times 10^{-4}$ 0.0246
For the statistical performance of $\widehat{\theta}$, this is tabulated in Table 2, which contains the bias (bias), median bias (mbias), standard deviation (std), bootstrap standard error (b-se), mean squared error (mse), and the scaled interquartile range (iqr) of the estimator. In particular, we use the nonparametric bootstrap to estimate the bootstrap standard error. We do this by resampling each simulated dataset with replacement and re-do the estimation procedure $99$ times. And we calculate iqr by dividing the interquartile range of our studentized estimates by 1.349.
The results give evidence that our $\widehat{\theta}$ is a consistent estimator for $\theta_{0}$, as the bias and standard deviation, and subsequently mean squared error, are decreasing with $L$. The reported iqr being close to $1$ indicates that our estimator has a normal-like tail behavior. The bootstrap standard error also approximates the standard deviation well. These comments apply to all bandwidths considered and the performances across bandwidths are comparable. Our simulation study thus supports our theoretical results and the recommendation that inference on the risk parameter can be done by using the nonparametric bootstrap.
Table 2: Estimation results for $\widehat{\theta}$.
$\theta_{true}$ $h$ $L$ bias mbias std b-se mse iqr
0 $sL^{-1/5}$ 250 0.1675 0.1881 0.2969 0.2789 0.1161 0.9857
500 0.1063 0.1121 0.2193 0.2163 0.0593 0.9668
1000 0.0686 0.0751 0.1672 0.1632 0.0326 0.9976
$sL^{-1/6}$ 250 0.1547 0.1811 0.2910 0.2731 0.1086 0.9764
500 0.1016 0.1041 0.2123 0.2100 0.0553 0.9459
1000 0.0706 0.0776 0.1615 0.1575 0.0311 0.9934
$sL^{-1/7}$ 250 0.1501 0.1738 0.2859 0.2683 0.1042 0.9722
500 0.0990 0.1037 0.2044 0.2046 0.0515 0.9520
1000 0.0728 0.0776 0.1558 0.1522 0.0296 0.9944
$0.5$ $sL^{-1/5}$ 250 0.0994 0.1182 0.2781 0.2728 0.0871 0.9894
500 0.0588 0.0535 0.1994 0.2043 0.0432 0.9766
1000 0.0376 0.0360 0.1504 0.1508 0.0240 0.9888
$sL^{-1/6}$ 250 0.0979 0.1194 0.2701 0.2646 0.0825 0.9812
500 0.0610 0.0577 0.1929 0.1970 0.0409 0.9881
1000 0.0444 0.0459 0.1441 0.1450 0.0227 0.9904
$sL^{-1/7}$ 250 0.1007 0.1177 0.2638 0.2581 0.0797 0.9731
500 0.0640 0.0614 0.1852 0.1907 0.0384 0.9638
1000 0.0496 0.0495 0.1384 0.1394 0.0216 1.0292
11 $sL^{-1/5}$ 250 0.0487 0.0547 0.2782 0.2918 0.0797 0.9706
500 0.0186 0.0107 0.1934 0.2093 0.0377 0.9834
1000 0.0153 0.0154 0.1440 0.1516 0.0210 0.9727
$sL^{-1/6}$ 250 0.0559 0.0674 0.2680 0.2792 0.0749 0.9753
500 0.0269 0.0205 0.1879 0.1997 0.0360 0.9700
1000 0.0249 0.0195 0.1378 0.1441 0.0196 0.9963
$sL^{-1/7}$ 250 0.0636 0.0728 0.2614 0.2693 0.0723 0.9723
500 0.0348 0.0287 0.1818 0.1917 0.0342 0.9397
1000 0.0315 0.0266 0.1327 0.1378 0.0186 1.0023
6 Empirical illustration
This section applies our estimator to real estate auction data from São Paulo. Real estate auctions constitute a large and active market in Brazil. Our sample consists of foreclosure apartments that we webscrapped from the website of a single large auctioneer, Zukerman (https://www.zukerman.com.br), which is recognized as the largest real estate auction platform in Brazil. The sellers are typically private and public banks, private companies that provide funds for borrowers, and the Court of Justice of the State of São Paulo (Tribunal de Justiça do Estado de São Paulo, TJ-SP).
Properties can be auctioned off for several legal reasons: (i) default on mortgage payments for more than six months; (ii) default on condominium maintenance fees; (iii) labor-related lawsuits; and (iv) other unpaid debt obligations. Auctions in categories (i) and (ii) are the most prevalent. Type (i) cases are classified as extrajudicial because the auction does not require judicial, or court, approval and are most often initiated by financial institutions when a property serving as collateral under fiduciary alienation is repossessed after borrower default. All other types generally require authorization by a court and typically consider the cases of unpaid loans, bankruptcy, or overdue condominium fees.
Our application involves auctions that may proceed to a second round if the property is not sold initially, in which case a new auction is held with a reduced reserve price. In the latter cases, judicial auctions typically apply a 50% discount on the reserve price when a property is not sold in the first round and extrajudicial auctions often set the second-round reserve price to the outstanding debt. To apply our methodology, we always take the reserve price in the first round to be the seller’s optimal reserve price. This is motivated by the fact that the first reserve is set based on actual appraisals of the property value, while the discounted reserve price in the second round is guided by either a judicial mandate or the property debt.
All sales follow the English auction format, in which bids are placed electronically in ascending order. The auction details including property information and auction schedule are publically available. The auction platform publishes all submitted bids in real time. We note that the second rounds, on average, take place 20 days after the first round. Our sample consists only of auctions with at least two bidders. We assume that all potential bidders submitted bids, which means we observe $N$ that corresponds to the number of different bidders in each auction. [^9] The assumptions on observing $N$ is a common one in practice, although models with endogenous bidders’ entry or unknown number of bidders have been studied in the literature.
The data we use contains 754 observations covering the period from 2017 to 2023, with the majority corresponding to auctions of type (i) and (ii). We did not collect data for cases of type (iii), as labor courts manage these auctions separately under distinct rules. The sample also excluded auctions deemed as outliers in terms of the reserve price, which we define as having the ratio of the reserve price to the size of the apartment in each auction that is larger than the 99th percentile or smaller than the 1st percentile of the sample. The price of the apartment is in Brazillian reais, and we measure it in R\$$100,000$s. [^10] All prices are expressed in constant January 2017 R\$. We use the size of the apartment (total area measured in sqm) as the covariate, $X$. To compute $W$, i.e. the value of the seller’s outside option, we subtract from the evaluation value of the property the debt registered in its sale report. We only have these information for 341 properties of the total sample. Since we do not use these information to estimate the quantile function of bidder’s private value, the quantiles are estimated using the whole sample. We then use the subsample with evaluation and debt values to estimate the seller’s risk parameter. Throughout this exercise, we assume the seller has a CRRA utility function.
Our estimation procedure then takes place over two stages. First, we estimate the bidder’s private value quantile function for the property. In a second step, we evaluate the seller’s CRRA risk parameter. We estimate $V(\alpha|X)$ using the AQR method for $\alpha=0.01,0.02,...,0.99$. [^11] In a very small proportion of auctions, the observed reserve price is larger than $\widehat{V}(0.99|X)$ or smaller than $\widehat{V}(0.01|X)$. These auctions are removed when estimating $\theta$ and during the model fit analysis. We consider three different bandwidth in our estimation: $h=0.1$, $h=0.15$, and $h=0.2$. The results show that our estimation methodology is robust for different bandwidth choices. To derive the standard error and some quantile levels of $\widehat{\theta}$, we use the nonparametric bootstrap. The bootstrap size is set at 99.
6.1 Descriptive statistics
We provide some descriptive statistics of our data. The variables involved are the reserve price ($R$), the winning bid ($B$), the size of the property ($X$), the outside option value ($W$), and the number of bidders in each auction ($N$). The summary statistics of the data containing the mean, standard deviation, and the quartiles are shown in Table 3.
Table 3: Descriptive statistics of the real estate data.
variable mean median 25-th pct 75-th pct std observations
$R$ 5.2990 3.5952 2.2305 6.1361 5.7411 754
$B$ 3.9331 2.7178 1.6736 4.7758 4.0066 754
$W$ 3.8883 2.5513 1.4385 4.4816 4.5412 341
$X$ 146.48 113.09 83.40 165.56 109.41 754
NN 5.8753 5 3 8 4.1901 754
Note that the reserve prices in the data tend to be higher than the winning bids, which is due to sales occurring in the second auction with lowered reserves. Observing bids below the reserve facilitates identification of the bidder’s valuations in the same manner to auctions with a secret or hidden reserve price (Elyakime et al., (1994), Andreyanov and Caoui, (2022)).
We provide the scatter plots of $X$ against $B$ and $R$ in Figure 2. The plots indicate a general trend that both winning bid and reserve price increase with the size of the property.
Figure 2: Scatter figures of $X$, $B$, and $R$.
6.2 Estimation results
We start by presenting figures of the estimates of the private value quantile function conditioning for property size at the three quartiles for different bandwidths.
Figure 3: Plots of $\widehat{V}(\cdot|X)$.
From the figures, we can see that $\widehat{V}(\alpha|X)$ is increasing in $\alpha$, slightly concave for small $\alpha$, and convex for large $\alpha$ for different $X$’s. The results using different $h$ are similar.
The estimation results for the risk-aversion parameter are given in Table 4, containing the bootstrap standard errors and the 2.5th and 97.5th percentiles.
Table 4: Results on $\widehat{\theta}$.
bandwidth $\widehat{\theta}$ b-se 2.5-th pct 97.5-th pct
$h=0.1$ 1.6025 0.2284 1.2373 2.0907
$h=0.15$ 1.5399 0.2538 1.1994 2.1625
$h=0.2$ 1.5242 0.2549 1.1387 2.0073
The results are qualitatively the same for all bandwidths. The $95\%$ bootstrapped coverage does not contain zero and the studentized statistic rejects the risk neutrality assumption in favor of risk aversion at any reasonable significant level.
6.3 Model fit
It is also instructive to consider the model fit under risk aversion. To do this, we simulate the winning bid and estimate the optimal reserve price using our estimated parameters and compute their CDFs to compare with the observed data. For the winning bid, we use the estimated $\widehat{V}(\alpha|X)$ to simulate the winning bid for each observed pair of ($N$, $X$) 1000 times, then combine the simulated winning bids across all different pairs of ($N,X$) in the sample to get a simulated CDF. To obtain the sample CDF, we use the empirical distribution of the observed winning bid data in the sample. The following figures show the comparison between the sample winning bid distribution and the simulated winning bid distribution using different bandwidths:
Figure 4: Model fit of winning bid distribution.
Visually, we can see that the simulated winning bid distributions fit the sample winning bid distributions very well for all bandwidths. This is complemented by the statistics in Table 5, which contains the bias (mean of simulated winning bids minus mean of sample winning bids), the percentage bias (bias divided by mean of sample winning bids), and the IMSE which is defined by
$IMSE_{B}=\int_{supp(B)}\left[F_{B}(x)-\widehat{F}_{B}(x)\right]^{2}dx,$
where $F_{B}(\cdot)$ is the CDF of the sample winning bids and $\widehat{F}_{B}(\cdot)$ is the CDF of the simulated winning bids. Table 5 shows that the percentage bias and the IMSE are all small, which suggest the fit of winning bid distribution is very good.
Table 5: Model fit of $\widehat{F}_{B}$.
$h$ bias percentage bias IMSEB
0.1 -0.0192 -0.49% 0.0045
0.15 -0.0192 -0.49% 0.0044
0.2 -0.0137 -0.35% 0.0045
We can also construct the model implied distribution of the reserve price. We use $\widehat{V}(\alpha|X)$, $\widehat{\theta}$, and the observed ($N,X,W$) to calculate the seller’s expected utility for $\alpha=0.01,0.02,...,0.99$, then find the optimal screening level $\alpha_{R}$ as the one that maximizes the seller’s expected utility. Finally, we use $\widehat{V}(\alpha_{R}|X)$ as the estimated optimal reserve price. Similar to the description above, we compare the CDF of observed reserve price and the CDF of estimated reserve price. The results are shown in the Figure 5.
Figure 5: Model fit of reserve price distribution.
Table 6 gives the bias, percentage bias, and IMSE of the estimated reserve price distribution, where the definition of bias, percentage bias, and IMSE is similar to that for the winning bid.
Table 6: Model fit of $\widehat{F}_{R}$.
$h$ bias percentage bias IMSER
0.1 0.0589 1.23% 0.0038
0.15 0.0556 1.18% 0.0039
0.2 0.0760 1.61% 0.0057
The figures and table presented above suggest our model generates the reserve price that fits the data well. Importantly, the results are qualitatively the same and are stable for all bandwidths considered.
As a simple counterfactual exercise, we construct CDFs of the reserve price distribution for a risk-neutral seller, which were constructed analogously as those in Figure 5. We provide these in Figure 6.
Figure 6: Counterfactual reserve price distribution.
We see that the distribution of the observed reserve price is almost first order stochastic dominated by the distribution of the counterfactual reserve price when the seller is risk neutral, which implies the risk-neutral seller’s reserve price tend to be higher than the risk-averse counterpart. To provide some quantitative comparisons, we calculate the amount of increase and the percentage increase in the counterfactual reserve price for the whole sample and for three subsamples around each quartile of the apartment size: small group (with area between $20$th and $30$th percentiles), medium group (with area between $45$th and $55$th percentiles), and large group (with area between $70$th and $80$th percentiles). The results are summarized in Table 7.
Table 7: Counterfactual reserve price.
$h$ sample increase in $R$ percentage increase
$0.1$ overall 0.6852 14.36%
small 0.6609 26.19%
medium 0.7251 20.61%
large 0.7498 13.23%
$0.15$ overall 0.6581 13.93%
small 0.6704 27.73%
medium 0.7001 20.00%
large 0.6432 11.40%
$0.2$ overall 0.6690 14.16%
small 0.6854 28.35%
medium 0.6625 18.92%
large 0.6533 11.58%
Overall, the seller’s reserve price would increase by 13% to 15% under the risk neutrality assumption, confirming that the seller’s optimal reserve price is decreasing in the degree of risk aversion.
7 Conclusion
In this paper we propose a framework to identify and estimate the seller’s risk aversion in ascending auctions. The conditions we use for identification are mild and are analogous to those assumed in the theoretical literature. On the estimation front, we apply the AQR approach of GG22 to estimate valuation quantiles of bidders from the winning bids and to estimate the risk parameter as a two-step semiparametric estimator. Our estimators perform well in a simulation study. We then apply our methodology to study the real estate auction data in Brazil, where we find statistical evidence that the sellers are risk-averse.
While we motivate our interest in seller’s risk aversion with the empirical observation of low reserve prices, it should be mentioned that there are other explanations presented in the theoretical literature to rationalize this. Other possibilities include endogenous entry (Levin and Smith, (1994)), affiliated types (Levin and Smith, (1996)), and risk averse bidders with interdependent values (Hu et al., (2019)). Our paper does not attempt to distinguish between these possibilities. In fact, to date, we are not aware of any works that have attempted this empirically, which would be an interesting direction of research. Nevertheless, we emphasize that allowing sellers to be risk-averse is an important advancement in empirical auctions as some counterfactuals and theoretical results require correct specification of the seller’s risk preference. Currently, empirical studies are assuming sellers are risk-neutral when performing counterfactual studies. Our paper shows how to accommodate risk-averse sellers in an ascending auction.
Since bidders’ behavior in a second-price auction is strategically equivalent to that of an ascending auction under IPV, our estimation strategy in this paper applies to second-price auctions. We believe our approach to model and estimate the seller’s risk aversion can be applied to other types of auctions as well. Beyond the auction format, other interesting aspects of our paper could be explored further. Firstly, we consider the traditional IPV paradigm. Various extensions from this framework have been proposed in the empirical auction literature, such as unobserved heterogeneity (Krasnokutskaya, (2011), Hu et al., (2013), and Luo and Xiao, (2023)), endogenous entry (Marmer et al., (2013), Gentry and Li, (2014), Chen et al., (2025)), and interdependent values (Gimenes and Guerre, (2020)). Even within the narrower domain of AQR applications, which have thus far been applied to ascending and first-price auctions, some practical aspects will benefit from further studies. For instance, while convergence rates for AQR estimators and their corresponding optimal bandwidths are relatively straightforwardly to derive, neither our paper nor GG22 provide practical guidance on bandwidth selection, and both rely on resampling methods for inference without formal justification. Despite encouraging finite-sample performance in Monte Carlo experiments and stable estimates with real data across bandwidth choices, further studies on these aspects will be useful for applied research.
Appendix
The appendix is organized into three subappendices. Appendix A gives the proof of Proposition 2; Appendix B gives the proofs of results in Section 3; Appendix C gives the proofs of results in Section 4.
A. Characterization of the optimal reserve price
In this proof, we omit $X$ for notational simplicity as the argument holds conditionally on $X$.
Proof of Proposition 2.
For part (a), omitting the auction covariates, equation (2) simplifies to
$\widetilde{\Pi}(r,w)=U\left(w\right)F\left(r\right)^{I}+IU\left(r\right)F\left(r\right)^{I-1}\left(1-F\left(r\right)\right)+I(I-1)\int_{r}^{\overline{v}}U\left(t\right)F(t)^{I-2}\left(1-F(t)\right)f(t)dt,$
for any $r$ and $w$. Taking partial derivative with respect to $r$ gives,
$\frac{\partial}{\partial r}\widetilde{\Pi}(r,w)=IF(r)^{I-1}f(r)\left\{U\left(w\right)+\frac{1}{f(r)}U^{\left(1\right)}(r)\left(1-F(r)\right)-U(r)\right\}.$
Let us define,
$h(r,w)=U(w)+\frac{1}{f(r)}U^{\left(1\right)}(r)\left(1-F(r)\right)-U(r).$ (15)
By M2(i), $f(r)>0$, thus the sign of $\frac{\partial}{\partial r}\widetilde{\Pi}(r,w)$ is determined by the sign of $h(r,w)$. Taking partial derivative with respect to $r$ gives
$\frac{\partial}{\partial r}h(r,w)=-2U^{\left(1\right)}(r)-U^{\left(1\right)}(r)\frac{f^{\left(1\right)}(r)\left(1-F(r)\right)}{f(r)^{2}}+U^{\left(2\right)}(r)\frac{1-F(r)}{f(r)}.$
By M2(iv), $U^{\left(2\right)}(\cdot)\leq 0$, thus we have
$\frac{\partial}{\partial r}h(r,w)\leq-2U^{\left(1\right)}(r)-U^{\left(1\right)}(r)\frac{f^{\left(1\right)}(r)\left(1-F(r)\right)}{f(r)^{2}}=-U^{\left(1\right)}(r)\left(2+\frac{f^{\left(1\right)}(r)\left(1-F(r)\right)}{f(r)^{2}}\right).$
By M2(ii), for all $v\in[\underline{v},\bar{v}]$,
$J^{\left(1\right)}(v)=1-\frac{-f(v)^{2}-\left(1-F(v)\right)f(v)}{f(v)^{2}}=2+\frac{\left(1-F(v)\right)f(v)}{f(v)^{2}}>0.$
Since $U^{\left(1\right)}(\cdot)>0$, there must be $\frac{\partial}{\partial r}h(r,w)\leq-U^{\left(1\right)}(r)J^{\left(1\right)}(r)<0$, which implies $h(\cdot,w)$ is strictly decreasing in $\underline{v},\bar{v}]$. Next, we check the boundary. Since $w\in[\underline{v},\bar{v})$,
$h(\underline{v},w)$ $=U(w)+\frac{1}{f(\underline{v})}U^{\prime}(\underline{v})-U(\underline{v})>0,$
$h(\bar{v},w)$ $=U(w)-U(\bar{v})<0.$
Therefore, there exists a unique $r\in(\underline{v},\bar{v})$ that maximizes $\Pi(r,w)$.
Now we consider part (b). As already shown, the optimal reserve price $r^{\ast}$ is the unique solution to $h(r^{\ast},w)=0$ and $\frac{\partial}{\partial r}h(r,w)<0$ for all $r\in(\underline{v},\bar{v})$ and $w$. By the Implicit Function Theorem, the optimal reserve price $r^{\ast}$ can be written as a map $w\mapsto r^{\ast}\left(w\right)$. It follows that $r^{\ast}\left(w\right)$ strictly increases with $w$, since
$\frac{dr^{\ast}\left(w\right)}{dw}=-\frac{\partial h(r,w)/\partial w}{\partial h(r,w)/\partial r},$
and $\frac{\partial}{\partial w}h(r,w)=U^{\left(1\right)}(w)>0$. To show $r^{\ast}$ strictly decreases with the seller’s risk aversion, consider $U_{1}(\cdot)$ and $U_{2}(\cdot)$ where there exists a strictly increasing and strictly concave function $\zeta(\cdot)$ such that $U_{2}(\cdot)=\zeta(U_{1}(\cdot))$. Denote the optimal reserve price for $U_{1}(\cdot)$ and $U_{2}(\cdot)$ by $r_{1}$ and $r_{2}$ respectively. By the result of part (a), $r_{1}$ and $r_{2}$ satisfy
$U_{1}(w)+\frac{1}{f(r_{1})}U_{1}^{(1)}(r_{1})(1-F(r_{1}))-U(r_{1})$ $=0,$
$U_{2}(w)+\frac{1}{f(r_{2})}U_{2}^{(1)}(r_{2})(1-F(r_{2}))-U(r_{2})$ $=0,$
which in turn implies
$U_{2}(r_{2})-\frac{1-F(r_{2})}{f(r_{2})}U_{2}^{\left(1\right)}(r_{2})=\zeta\left(U_{1}(r_{1})-\frac{1-F(r_{1})}{f(r_{1})}U_{1}^{\left(1\right)}(r_{1})\right).$
Since $\zeta(\cdot)$ is strictly concave,
$\zeta\left(U_{1}(r_{1})-\frac{1-F(r_{1})}{f(r_{1})}U_{1}^{(1)}(r_{1})\right)$ $<\zeta(U_{1}(r_{1}))-\zeta^{(1)}(U_{1}(r_{1}))\frac{1-F(r_{1})}{f(r_{1})}U_{1}^{(1)}(r_{1})$
$=U_{2}(r_{1})-\frac{1-F(r_{1})}{f(r_{1})}U_{2}^{(1)}(r_{1}).$
Therefore, there must be
$U_{2}(r_{2})-\frac{1-F(r_{2})}{f(r_{2})}U_{2}^{\left(1\right)}(r_{2})<U_{2}(r_{1})-\frac{1-F(r_{1})}{f(r_{1})}U_{2}^{\left(1\right)}(r_{1}).$
As shown in part (a), we know $U_{2}(r)-\frac{1-F(r)}{f(r)}U_{2}^{\left(1\right)}(r)$ is strictly increasing, thus $r_{2}<r_{1}$$\blacksquare$
B. Asymptotic Theory for Quantile Function Estimators
[…]
Before getting into the technical details, it is instructive to clarify how our objects of interest differ from those in GG. In a first-price auction, the quantile function of a bidder’s valuation depends on $B\left(\cdot\right)$ and $B^{\left(1\right)}\left(\cdot\right)$ – the quantile of an individual’s bid and its derivative. GG22 show that the estimation error of $B\left(\cdot\right)$ converges to zero at a faster rate than the counterparts of the estimator of $B^{\left(1\right)}\left(\cdot\right)$, so they derive leading bias and variance expressions for the quantile derivative estimator of the observed bids, but only provide a convergence rate for the quantile estimator itself. In contrast, in an ascending auction, only the quantile of the winning bids is needed to recover the quantile of the bidder’s valuation. Our results provide bias and variance expressions for the quantile estimator in this setting.
Our quantile estimation setup below is a simplification of GG22’s general framework. We use what they call Augmented Quantile Regression (AQR), which is presented in their main text. In contrast, their appendix provides proofs for Augmented Sieve Quantile Regression (ASQR), where the linear quantile specification includes an increasing number of additive terms, leading to a nonparametric specification. As a result, our appendix is much lighter in notation and the proofs simplified.
Our presentation in this sub-appendix aims to be as concise as possible. We outline key arguments, omit steps that follow directly from GG22, refer readers to the original source, and focus only on the elements that require modification from their work. This approach avoids unnecessary repetition of GG’s appendix, which is a diligently crafted piece spanning several sub-appendices that would remain lengthy even without ASQR components.
In what follows, we refer to GG22’s supplementary material as GG-SM. This document can found online at https://doi.org/10.1016/j.jeconom.2021.02.009. The relevant sections are most of Appendices B, C, D, and E. To facilitate readers, we try to use the same notation as in GG22 when possible.
B.1 Outline
We start by defining the sample and population objective functions with a reparameterized parameter. As standard in the asymptotic analysis local polynomial regression, for example see Fan and Gijbels (1996), we rescale the parameters to avoid degeneracy of the “design” matrix, $E\left[P\left(X_{l},th\right)P\left(X_{l},th\right)^{\top}\right]$, as $h\rightarrow 0$. We do this by changing the parameter to $\mathsf{b}=Hb$, where $H=diag\left(1,\ldots,h^{s}\right)\otimes\mathrm{I}_{D+1}$, so that $P\left(X_{l},th\right)^{\top}b=P\left(X_{l},t\right)^{\top}\mathsf{b}$.
Then, we define $\widehat{\mathsf{b}}\left(\alpha\right)=\arg\min_{b}\widehat{\mathsf{R}}\left(\mathsf{b};\alpha\right)$, where
$\widehat{\mathsf{R}}\left(\mathsf{b};\alpha\right)=\frac{1}{L}\sum\limits_{l=1}^{L}\int_{-\frac{\alpha}{h}}^{\frac{1-\alpha}{h}}\rho_{\alpha+ht}\left(B_{l}-P\left(X_{l},t\right)^{\top}\mathsf{b}\right)K\left(t\right)dt,$ (16)
so that $\widehat{b}\left(\alpha\right)=H^{-1}\widehat{\mathsf{b}}\left(\alpha\right)$. We use $b\left(\alpha\right)$ to denote the vector of true coefficients of the quantile slopes and their derivatives, $b\left(\alpha\right)=\left[\beta\left(\alpha\right)^{\top},\beta^{\left(1\right)}\left(\alpha\right)^{\top},\ldots,\beta^{\left(s+1\right)}\left(\alpha\right)^{\top}\right]^{\top}$ and define $\mathsf{b}\left(\alpha\right)=Hb\left(\alpha\right)$. Since we are interested in $B\left(\alpha|X\right)=X_{1}^{\top}\beta\left(\alpha\right)$, and later on its derivatives, it will be useful to define row vectors of size $\left(s+1\right)$, denoted by $S_{j}$, so that of $S_{j}\$has 0 in each of its entry other than $1$ in the $\left(j+1\right)$-th entry for $j=0,1,\ldots,s$. For example, $S_{0}=\left[1,\ldots,0\right]$, $S_{1}=\left[0,1,\ldots,0\right]$, and so on. Then let $\mathsf{S}_{j}=S_{j}\otimes\mathrm{I}_{D+1}$, so we can estimate $B\left(\alpha|x\right)$ by $\widehat{B}\left(\alpha|x\right)=x_{1}^{\top}\mathsf{S}_{0}\widehat{\mathsf{b}}\left(\alpha\right)$ and $\widehat{B}^{\left(j\right)}\left(\alpha|x\right)=x_{1}^{\top}\mathsf{S}_{j}\widehat{\mathsf{b}}\left(\alpha\right)/h^{j}$ for $j=1,\ldots,s$.
The properties of $\widehat{\mathsf{b}}\left(\alpha\right)$ can be conveniently analyzed using its Bahadur representation:
$\widehat{\mathsf{b}}\left(\alpha\right)=\overline{\mathsf{b}}\left(\alpha\right)+\widehat{\mathsf{e}}\left(\alpha\right)+\widehat{\mathsf{d}}\left(\alpha\right),$ (17)
where $\overline{\mathsf{b}}\left(\alpha\right)=\arg\min_{\mathsf{b}}\overline{\mathsf{R}}\left(\mathsf{b};\alpha\right)$, $\widehat{\mathsf{e}}\left(\alpha\right)=-\left[\overline{\mathsf{R}}^{\left(2\right)}\left(\overline{\mathsf{b}}\left(\alpha\right);\alpha\right)\right]^{-1}\widehat{\mathsf{R}}^{\left(1\right)}\left(\overline{\mathsf{b}}\left(\alpha\right);\alpha\right)$ with $\overline{\mathsf{R}}^{\left(2\right)}\left(\overline{\mathsf{b}}\left(\alpha\right);\alpha\right)=E\left[\widehat{\mathsf{R}}^{\left(2\right)}\left(\overline{\mathsf{b}}\left(\alpha\right);\alpha\right)\right]$ that can be shown to exist and have full rank for small enough hh [^12] Lemma B.3(i) shows in GG-SM shows $\widehat{\mathsf{R}}\left(\mathsf{b};\alpha\right)$ and $E\left[\widehat{\mathsf{R}}\left(\mathsf{b};\alpha\right)\right]$ are twice continuously differentiable for $\mathsf{b}\in\mathcal{B}_{\infty}\left(\mathsf{b}\left(\alpha\right),C_{0}h\right)$ for some $C_{0}$ and small enough $h$. Moreover, $\overline{\mathsf{R}}^{\left(2\right)}\left(\mathsf{b}\left(\alpha\right);\alpha\right)$ has full rank for all $\alpha\in\left[0,1\right]$ by Lemma B.3(ii). This is relevant since $\left\|\overline{\mathsf{b}}\left(\alpha\right)-\mathsf{b}\left(\alpha\right)\right\|=o\left(h^{s+1}\right)$, see equation (C.2) which is shown in the proof of Theorem C.3, and Lemma B.3(i) shows $\sup_{\alpha\in\left[0,1\right]\mathsf{b}^{1},\mathsf{b}^{0}\in\mathcal{B}_{\infty}\left(\mathsf{b}\left(\alpha\right),C_{0}h\right)}\frac{\left\|\overline{\mathsf{R}}^{\left(2\right)}\left(\mathsf{b}^{1};\alpha\right)-\overline{\mathsf{R}}^{\left(2\right)}\left(\mathsf{b}^{0};\alpha\right)\right\|}{\left\|\mathsf{b}^{1}-\mathsf{b}^{0}\right\|/\left(\alpha\left(1-\alpha\right)+h\right)}=O\left(1\right)$ for some $C_{0}$ and small enough $h$., and $\widehat{\mathsf{d}}\left(\alpha\right)=\widehat{\mathsf{b}}\left(\alpha\right)-\overline{\mathsf{b}}\left(\alpha\right)-\widehat{\mathsf{e}}\left(\alpha\right)$.
Let $\widehat{\mathsf{b}}^{\ast}\left(\alpha\right)$ denote $\overline{\mathsf{b}}\left(\alpha\right)+\widehat{\mathsf{e}}\left(\alpha\right)$. $\widehat{\mathsf{b}}^{\ast}\left(\alpha\right)$ is the unique minimizer of the following quadratic approximation of $\widehat{\mathsf{R}}\left(\mathsf{b};\alpha\right)$:
$\widehat{\mathsf{R}}\left(\overline{\mathsf{b}}\left(\alpha\right);\alpha\right)+\left(\mathsf{b}-\overline{\mathsf{b}}\left(\alpha\right)\right)^{\top}\widehat{\mathsf{R}}^{\left(1\right)}\left(\overline{\mathsf{b}}\left(\alpha\right);\alpha\right)+\frac{1}{2}\left(\mathsf{b}-\overline{\mathsf{b}}\left(\alpha\right)\right)^{\top}\overline{\mathsf{R}}^{\left(2\right)}\left(\overline{\mathsf{b}}\left(\alpha\right);\alpha\right)\left(\mathsf{b}-\overline{\mathsf{b}}\left(\alpha\right)\right).$
The validity of the quadratic approximation will be confirmed by the convergence rate of $\widehat{\mathsf{d}}$.
In the AQR framework, $\overline{\mathsf{b}}\left(\alpha\right)$ can be interpreted as the pseudo-true parameter. Then $x_{1}^{\top}\mathsf{S}_{0}\left(\overline{\mathsf{b}}\left(\alpha\right)-\mathsf{b}\left(\alpha\right)\right)$ captures the polynomial approximation bias of $B\left(\alpha|x\right)$. It can be shown that $\widehat{\mathsf{e}}\left(\alpha\right)$ is a zero mean vector, and suitably scaled components of $\widehat{\mathsf{e}}\left(\alpha\right)$ satisfy the central limit theorem (CLT).
B.2 Proofs of results
We only provide the proofs of Lemmas 1–3. Propositions 3–5 immediately follow from these lemmas.
Proof of Lemma 1.
Consider the following decomposition:
$\widehat{B}\left(\alpha|x\right)-B\left(\alpha|x\right)$ $=$ $J_{B}\left(\alpha,x\right)+J_{S}\left(\alpha,x\right)+J_{R}\left(\alpha,x\right),\text{ where}$
$J_{B}\left(\alpha,x\right)$ $=$ $x_{1}^{\top}\mathsf{S}_{0}\left(\overline{\mathsf{b}}\left(\alpha\right)-\mathsf{b}\left(\alpha\right)\right),$ (18)
$J_{S}\left(\alpha,x\right)$ $=$ $x_{1}^{\top}\mathsf{S}_{0}\widehat{\mathsf{e}}\left(\alpha\right),$ (19)
$J_{R}\left(\alpha,x\right)$ $=$ $x_{1}^{\top}\mathsf{S}_{0}\widehat{\mathsf{d}}\left(\alpha\right),\$(20)
where $\left(J_{B}\left(\alpha,x\right),J_{S}\left(\alpha,x\right),J_{R}\left(\alpha,x\right)\right)$ are respectively the bias, leading stochastic term, and remainder term.
[…]
$\widehat{\mathsf{R}}^{\left(1\right)}\left(\mathsf{b};\alpha\right)$ $=$ $\frac{1}{L}\sum\limits_{l=1}^{L}\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\left(\mathbf{1}\left[B_{l}\leq P\left(X_{l},t\right)^{\top}\mathsf{b}\right]-\left(\alpha+ht\right)\right)P\left(X_{l},t\right)K\left(t\right)dt$
$E\left[\widehat{\mathsf{R}}^{\left(1\right)}\left(\mathsf{b};\alpha\right)\right]$ $=$ $\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\left(E\left[\mathbf{1}\left[B_{l}\leq P\left(X_{l},t\right)^{\top}\mathsf{b}\right]\right]-\left(\alpha+ht\right)\right)P\left(X_{l},t\right)K\left(t\right)dt$
$=$ $\int\left(\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\left(G\left(P\left(x,t\right)^{\top}\mathsf{b}|x\right)-\left(\alpha+ht\right)\right)P\left(x,t\right)K\left(t\right)dt\right)f\left(x\right)dx,$
where $f\left(\cdot\right)$ is the PDF of $X_{l}$, $G\left(\cdot\right)$ is the conditional CDF of $B_{l}$ given $X_{l}$, $\underline{t}_{\alpha,h}=-\min\left(1,\frac{\alpha}{h}\right)$ and $\overline{t}_{\alpha,h}=\max\left(1,\frac{1-\alpha}{h}\right)$, noting that $\mathcal{T}_{\alpha,h}=\left[\underline{t}_{\alpha,h},\overline{t}_{\alpha,h}\right]$ serve as the effective range of integration since $K\left(\cdot\right)$ is truncated.
GG-SM prove, in Step 1 of the proof of their Theorem C.3, that there exists $\overline{\mathsf{b}}\left(\alpha\right)\inB$$_{\infty}\left(\mathsf{b}\left(\alpha\right),C_{0}h\right)$ for some $C_{0}$ and small enough $h$, that $E\left[\widehat{\mathsf{R}}^{\left(1\right)}\left(\overline{\mathsf{b}}\left(\alpha\right);\alpha\right)\right]=0$. Let $\Psi\left(t|x,\mathsf{b}\right)$ denote $P\left(X_{l},t\right)^{\top}\mathsf{b}$ for $t\in\mathcal{T}_{\alpha,h}$, which can be shown to be invertible when $\mathsf{b}$ is in a vicinity of $\mathsf{b}\left(\alpha\right)$, which applies here for $\mathsf{b}\inB$$_{\infty}\left(\mathsf{b}\left(\alpha\right),C_{0}h\right)$. Expanding the first-order condition, $E\left[\widehat{\mathsf{R}}^{\left(1\right)}\left(\overline{\mathsf{b}}\left(\alpha\right);\alpha\right)\right]=0$, and follow the arguments in Step 2 in the proof of Theorem C.3 in GG-SM yields:
$0$ $=$ $\int\left(\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\overline{g}\left(\alpha|t,x\right)\left(\Psi\left(t|x,\overline{\mathsf{b}}\left(\alpha\right)\right)-\Psi\left(t|x,\mathsf{b}\left(\alpha\right)\right)\right)P\left(x,t\right)K\left(t\right)dt\right)f\left(x\right)dx$
$+\int\left(\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\overline{g}\left(\alpha|t,x\right)\left(\Psi\left(t|x,\mathsf{b}\left(\alpha\right)\right)-B\left(\alpha+ht\right)\right)P\left(x,t\right)K\left(t\right)dt\right)f\left(x\right)dx,$
where $\overline{g}\left(\alpha|t,x\right)=\int_{0}^{1}g\left(\Psi\left(t|x,\overline{\mathsf{b}}\left(\alpha\right)\right)+u\left(B\left(\alpha+ht|x\right)-\Psi\left(t|x,\mathsf{b}\left(\alpha\right)\right)\right)\right)du$. Applying Taylor’s expansion in the second term, viz., parts (iii) and (iv) in Lemma B.2 of GG-SM, gives
$\mathsf{\check{R}}^{\left(2\right)}\left(\alpha\right)\left(\overline{\mathsf{b}}\left(\alpha\right)-\mathsf{b}\left(\alpha\right)\right)$ $=$ $\mathsf{\check{R}}^{\left(2\right)}\left(\alpha\right)h^{s+1}\mathsf{\check{b}}_{s+1}\left(\alpha\right)+o\left(h^{s+1}\right),\text{ where}$ (21)
$\mathsf{\check{b}}_{s+1}\left(\alpha\right)$ $=$ $\left[\mathsf{\check{R}}^{\left(2\right)}\left(\alpha\right)\right]^{-1}\int\left(\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\overline{g}\left(\alpha|t,x\right)\frac{t^{s+1}B^{\left(s+1\right)}\left(\alpha|x\right)}{\left(s+1\right)!}P\left(x,t\right)K\left(t\right)dt\right)f\left(x\right)dx,$
$\mathsf{\check{R}}^{\left(2\right)}\left(\alpha\right)$ $=$ $\int\left(\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\overline{g}\left(\alpha|t,x\right)P\left(x,t\right)P\left(x,t\right)^{\top}K\left(t\right)dt\right)f\left(x\right)dx.$
Note that we perform a Taylor’s expansion upto the $\left(s+1\right)$-th term while GG-SM do it to the $\left(s+2\right)$-th term. This is because the quantile function of optimal first price auction bids has one more derivative than the quantile function of the underlying bidder’s valuation.
GG-SM show, as part of the proof of their Lemma C.2, that
$\max_{\left(\alpha,x\right)\in\left[0,1\right]\times\mathcal{X}}\max_{t\in\left[\underline{t}_{\alpha,h}+C_{1}\epsilon_{1},\overline{t}_{\alpha,h}-C_{1}\epsilon_{1}\right]}\left|\overline{g}\left(\alpha|t,x\right)-\frac{1}{B^{\left(1\right)}\left(\alpha|x\right)}\right|=O\left(h\right),$
for some positive constant $C_{1}$ and $\epsilon_{1}=o\left(h\right)$, and
$\max_{\left(\alpha,x\right)\in\left[0,1\right]\times\mathcal{X}}\left\|\mathsf{\check{R}}^{\left(2\right)}\left(\alpha\right)-\Omega_{h}\left(\alpha\right)\otimes\mathbf{P}_{0}\left(\alpha\right)\right\|=O\left(h\right),\text{ where}$
$\Omega_{h}\left(\alpha\right)=\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\pi\left(t\right)\pi\left(t\right)^{\top}K\left(t\right)dt\text{ and }\mathbf{P}_{0}\left(\alpha\right)=E\left[\frac{X_{l}X_{l}^{\top}}{B^{\left(1\right)}\left(\alpha|X_{l}\right)}\right].$
Therefore, since $h=o\left(1\right)$ and dominated convergence applies, we have:
$\mathsf{\check{b}}_{s+1}\left(\alpha\right)$
$=$ $\left[\Omega_{h}\left(\alpha\right)\otimes\mathbf{P}_{0}\left(\alpha\right)+o\left(1\right)\right]^{-1}\int\left(\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\left(\frac{t^{s+1}B^{\left(s+1\right)}\left(\alpha|x\right)}{\left(s+1\right)!B^{\left(1\right)}\left(\alpha|x\right)}+o\left(1\right)\right)P\left(x,t\right)K\left(t\right)dt\right)f\left(x\right)dx$
$=$ $\left[\Omega_{h}\left(\alpha\right)^{-1}\otimes\mathbf{P}_{0}\left(\alpha\right)^{-1}\right]\left[\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\frac{t^{s+1}\pi\left(t\right)}{\left(s+1\right)!}K\left(t\right)dt\otimes\int\frac{x_{1}x_{1}^{\top}}{B^{\left(1\right)}\left(\alpha|x\right)}f\left(x\right)dx\beta^{\left(s+1\right)}\left(\alpha\right)\right]+o\left(1\right)$
$=$ $\Omega_{h}\left(\alpha\right)^{-1}\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\frac{t^{s+1}\pi\left(t\right)}{\left(s+1\right)!}K\left(t\right)dt\otimes\beta^{\left(s+1\right)}\left(\alpha\right)+o\left(1\right).$
Since $\mathsf{S}_{0}\left[\Omega_{h}\left(\alpha\right)^{-1}\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\frac{t^{s+1}K\left(t\right)}{\left(s+1\right)!}\pi\left(t\right)dt\otimes\beta^{\left(s+1\right)}\left(\alpha\right)\right]=S_{0}\Omega_{h}\left(\alpha\right)^{-1}\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\frac{t^{s+1}K\left(t\right)}{\left(s+1\right)!}\pi\left(t\right)dt\beta^{\left(s+1\right)}\left(\alpha\right)$, the proof is completed as
$x_{1}^{\top}\mathsf{S}_{0}\left(\overline{\mathsf{b}}\left(\alpha\right)-\mathsf{b}\left(\alpha\right)\right)$ $=$ $h^{s+1}x_{1}^{\top}\beta^{\left(s+1\right)}\left(\alpha\right)S_{0}\Omega_{h}\left(\alpha\right)^{-1}\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\frac{t^{s+1}\pi\left(t\right)}{\left(s+1\right)!}K\left(t\right)dt+o\left(h^{s+1}\right)$
$=$ $h^{s+1}B^{\left(s+1\right)}\left(\alpha|x\right)S_{0}\Omega_{h}\left(\alpha\right)^{-1}\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\frac{t^{s+1}\pi\left(t\right)}{\left(s+1\right)!}K\left(t\right)dt+o\left(h^{s+1}\right).$
Next, we consider $J_{S}\left(\alpha,x\right)$. We start with the variance expression for $\widehat{\mathsf{e}}\left(\alpha\right)$ given in GG-SM. In particular, when proving their Lemma B.6 in Appendix F.3.3, they show that $Var\left(\widehat{\mathsf{e}}\left(\alpha\right)\right)=\left(\mathsf{V}_{e}\left(\alpha\right)+o\left(h\right)\right)/L$, [^13] There are a couple of typos in GG-SM related to $Var\left(\widehat{\mathsf{e}}\left(\alpha\right)\right)$ that are easy to verify. First, at the bottom of their page 134, they were meant to say $Var\left(\widehat{\mathsf{e}}\left(\alpha\right)\right)=\left(\mathsf{V}_{e}+o\left(h\right)\right)/L$ instead of $Var\left(\widehat{\mathsf{e}}\left(\alpha\right)\right)=\mathsf{V}_{e}/L+o\left(h\right)$. Second, the last component of $\mathsf{V}_{e}$ given in the second display of page 135 should be $S_{0}^{\top}S_{1}+S_{1}^{\top}S_{0}$ and not $S_{0}S_{1}^{\top}+S_{1}S_{0}^{\top}$. where
$\mathsf{V}_{e}\left(\alpha\right)$ $=$ $\alpha\left(1-\alpha\right)\left[S_{0}^{\top}S_{0}\right]\otimes\left[\mathbf{P}_{0}\left(\alpha\right)^{-1}\mathbf{PP}_{0}\left(\alpha\right)^{-1}\right]$
$+h\alpha\left(1-\alpha\right)\left[S_{1}^{\top}S_{0}\right]\otimes\left[\mathbf{P}_{0}\left(\alpha\right)^{-1}\mathbf{P}_{1}\mathbf{P}_{0}\left(\alpha\right)^{-1}\mathbf{PP}_{0}\left(\alpha\right)^{-1}\right]$
$+h\alpha\left(1-\alpha\right)\left[S_{0}^{\top}S_{1}\right]\otimes\left[\mathbf{P}_{0}\left(\alpha\right)^{-1}\mathbf{PP}_{0}\left(\alpha\right)^{-1}\mathbf{P}_{1}\mathbf{P}_{0}\left(\alpha\right)^{-1}\right]$
$+h\left[\Omega_{h}^{-1}\mathbf{\Pi}_{m}\Omega_{h}^{-1}-\left(S_{0}^{\top}S_{1}+S_{1}^{\top}S_{0}\right)\right]\otimes\left[\mathbf{P}_{0}\left(\alpha\right)^{-1}\mathbf{PP}_{0}\left(\alpha\right)^{-1}\right],$
$\mathbf{\Pi}_{m}\left(\alpha\right)$ $=$ $\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\min\left(t_{1},t_{2}\right)\pi\left(t\right)\pi\left(t\right)^{\top}K\left(t_{1}\right)K\left(t_{2}\right)dt_{1}dt_{2},$
$\mathbf{P}$ $=$ $E\left[X_{l}X_{l}^{\top}\right]\text{ and }\mathbf{P}_{0}\left(\alpha\right)=E\left[\frac{X_{l}X_{l}^{\top}}{B^{\left(1\right)}\left(\alpha|X_{l}\right)}\right].$
It follows that $Var\left(\mathsf{S}_{0}\widehat{\mathsf{e}}\left(\alpha\right)\right)=\alpha\left(1-\alpha\right)\mathbf{P}_{0}\left(\alpha\right)^{-1}\mathbf{PP}_{0}\left(\alpha\right)^{-1}/L+O\left(h/L\right)$, and we have $Var\left(\sqrt{L}\mathsf{S}_{0}\widehat{\mathsf{e}}\left(\alpha\right)\right)=\Sigma_{h}\left(\alpha\right)+o\left(1\right)$ as desired.
The central limit theorem (CLT) result follows from the same argument used in the proof of Theorem 3 (and A.3) in GG-SM that can be found in their Appendix E.2. Write $\left(\frac{L}{x_{1}^{\top}\Sigma_{h}\left(\alpha\right)x_{1}}\right)^{1/2}x_{1}^{\top}\mathsf{S}_{0}\widehat{\mathsf{e}}\left(\alpha\right)=\sum\limits_{l=1}^{L}r_{l}\left(\alpha|x\right)$, so that
$r_{l}\left(\alpha|x\right)$ $=$ $\left(\frac{1}{Lx_{1}^{\top}\Sigma_{h}\left(\alpha\right)x_{1}}\right)^{1/2}x_{1}^{\top}\mathsf{S}_{0}\left[\overline{\mathsf{R}}^{\left(2\right)}\left(\overline{\mathsf{b}}\left(\alpha\right);\alpha\right)\right]^{-1}$
$\times\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\left(\mathbf{1}\left[B_{l}\leq P\left(X_{l},t\right)^{\top}\overline{\mathsf{b}}\left(\alpha\right)\right]-\left(\alpha+ht\right)\right)P\left(X_{l},t\right)K\left(t\right)dt.$
Since $\left|Var\left(r_{l}\left(\alpha|x\right)\right)-1\right|=o\left(1\right)$, we can bound $\left|E\left[r_{l}^{3}\left(\alpha|x\right)\right]\right|$ by
$\left|r_{l}\left(\alpha|x\right)\right|Var\left(r_{l}\left(\alpha|x\right)\right)\leq O\left(\frac{1}{L}\right)=o\left(1\right),$
thus CLT applies.
For the remainder term, the proof that $\sqrt{L}J_{R}\left(\alpha,x\right)=o_{p}\left(1\right)$ follows immediately from applying Theorem D.1 in GG-SM, where they show that
$\sup_{\alpha\in\left[0,1\right]}\left\|\frac{Lh^{1/2}}{\left(h+\alpha\left(1-\alpha\right)\right)^{1/2}\log L}\widehat{\mathsf{d}}\left(\alpha\right)\right\|=O_{p}\left(1\right),$
under the condition $\frac{\log^{2}L}{Lh}=o\left(1\right)$.
For the uniform rate, the result will follow from the triangle inequality once we verify:
$\sup_{\left(\alpha,x\right)\in\left[0,1\right]\times\mathcal{X}}\left|J_{B}\left(\alpha,x\right)\right|$ $=$ $O\left(h^{s+1}\right),$
$\sup_{\left(\alpha,x\right)\in\left[0,1\right]\times\mathcal{X}}\left|J_{S}\left(\alpha,x\right)\right|$ $=$ $O_{p}\left(\sqrt{\frac{\log L}{L}}\right),$
$\sup_{\left(\alpha,x\right)\in\left[0,1\right]\times\mathcal{X}}\left|J_{R}\left(\alpha,x\right)\right|$ $=$ $o_{p}\left(\sqrt{\frac{\log L}{L}}\right).\$
Since $\mathcal{X}$ is compact, the pointwise bias rate holds uniformly as other components in $\mathsf{Bias}_{h}\left(\cdot\right)$ are bounded.
GG-SM have shown the required rate for $J_{S}\left(\cdot\right)$ in their Lemma B.6(ii).
Lastly, for $J_{R}\left(\cdot\right)$, this follows from
$\sup_{\alpha\in\left[0,1\right]}\left\|\widehat{\mathsf{d}}\left(\alpha\right)\right\|=O_{p}\left(\frac{\log L}{Lh^{1/2}}\right),$
[…]
Proof of Lemma 2. From the Bahadur representation, (17), we have:
$\widehat{B}^{\left(j\right)}\left(\alpha|x\right)-B^{\left(j\right)}\left(\alpha|x\right)=\frac{x_{1}^{\top}\mathsf{S}_{j}\left(\overline{\mathsf{b}}\left(\alpha\right)-\mathsf{b}\left(\alpha\right)\right)}{h^{j}}+\frac{x_{1}^{\top}\mathsf{S}_{j}\widehat{\mathsf{e}}\left(\alpha\right)}{h^{j}}+\frac{x_{1}^{\top}\mathsf{S}_{j}\widehat{\mathsf{d}}\left(\alpha\right)}{h^{j}}.\$
The three terms above respectively represent the bias, leading stochastic term, and remainder term. It suffices to show the following:
$\sup_{\left(\alpha,x\right)\in\left[0,1\right]\times\mathcal{X}}\left|\frac{x_{1}^{\top}\mathsf{S}_{j}\left(\overline{\mathsf{b}}\left(\alpha\right)-\mathsf{b}\left(\alpha\right)\right)}{h^{j}}\right|$ $=$ $O\left(h^{s+1-j}\right),$
$\sup_{\left(\alpha,x\right)\in\left[0,1\right]\times\mathcal{X}}\left|\frac{x_{1}^{\top}\mathsf{S}_{j}\widehat{\mathsf{e}}\left(\alpha\right)}{h^{j}}\right|$ $=$ $O_{p}\left(\sqrt{\frac{\log L}{Lh^{2j-1}}}\right),$
$\sup_{\left(\alpha,x\right)\in\left[0,1\right]\times\mathcal{X}}\left|\frac{x_{1}^{\top}\mathsf{S}_{j}\widehat{\mathsf{d}}\left(\alpha\right)}{h^{j}}\right|$ $=$ $o_{p}\left(\sqrt{\frac{\log L}{Lh^{2j-1}}}\right).$
The form of the bias can be obtained from the expansion of $\overline{\mathsf{b}}\left(\alpha\right)-\mathsf{b}\left(\alpha\right)=h^{s+1}\mathsf{\check{b}}_{s+1}\left(\alpha\right)+o\left(h^{s+1}\right)$, as shown in (21), where $\mathsf{\check{b}}_{s+1}\left(\alpha\right)$ is given in (B.2 Proofs of results). Then,
$x_{1}^{\top}\mathsf{S}_{j}\left(\overline{\mathsf{b}}\left(\alpha\right)-\mathsf{b}\left(\alpha\right)\right)=h^{s+1-j}x_{1}^{\top}\beta^{\left(s+1\right)}\left(\alpha\right)S_{j}\Omega_{h}\left(\alpha\right)^{-1}\int_{\underline{t}_{\alpha,h}}^{\overline{t}_{\alpha,h}}\frac{t^{s+1}\pi\left(t\right)}{\left(s+1\right)!}K\left(t\right)dt+o\left(h^{s+1-j}\right).$
It follows from the Inverse Function Theorem that $\beta^{\left(s+1\right)}\left(\cdot\right)$ is continuous on $\left[0,1\right]$, as $\gamma\left(\cdot\right)$, which has $s+1$ continuous derivatives, is a composite function of $\beta\left(\cdot\right)$ and a differentiable and strictly increasing function, $\phi\left(\cdot\right)$. Since $\mathcal{X}$ is compact, the bias is $O\left(h^{s+1-j}\right)$ uniformly.
We next consider the leading stochatic term. GG-SM prove this in their Lemma B.6 when $j=1$, by showing that
$\sup_{\left(\alpha,x\right)\in\left[0,1\right]\times\mathcal{X}}\left|\frac{x_{1}^{\top}\mathsf{S}_{j}\widehat{\mathsf{e}}\left(\alpha\right)}{h^{1/2}}\right|=O_{p}\left(\sqrt{\frac{\log L}{L}}\right),$
see their equation (F.3). The same proof in fact applies to cases when $j>1$ as long as $Var\left(\mathsf{S}_{j}\widehat{\mathsf{e}}\left(\alpha\right)/h^{1/2}\right)$ is uniformly bounded for $\alpha\in\left[0,1\right]$. This is indeed the case, as it can be seen from (B.2 Proofs of results) that:
$Var\left(\mathsf{S}_{j}\widehat{\mathsf{e}}\left(\alpha\right)\right)=h\left[S_{j}\Omega_{h}^{-1}\mathbf{\Pi}_{m}\Omega_{h}^{-1}S_{j}^{\top}\right]\otimes\left[\mathbf{P}_{0}\left(\alpha\right)^{-1}\mathbf{PP}_{0}\left(\alpha\right)^{-1}\right]\text{ for all }j>0;$
the expression follows from $S_{j}S_{j^{\prime}}^{\top}=0$ whenever $j^{\prime}\neq j$.
For the remainder term, we make use of of the fact that
$\sup_{\alpha\in\left[0,1\right]}\left\|\widehat{\mathsf{d}}\left(\alpha\right)\right\|_{\infty}=O_{p}\left(\frac{\log L}{Lh^{1/2}}\right),$
which was shown in the proof of Theorem D.1 in GG-SM – as implied by their equation (D.4). Then,
$\sup_{\left(\alpha,x\right)\in\left[0,1\right]\times\mathcal{X}}\left|\frac{x_{1}^{\top}\mathsf{S}_{j}\widehat{\mathsf{d}}\left(\alpha\right)}{h^{j}}\right|$ $=$ $O_{p}\left(\frac{\log L}{Lh^{j+1/2}}\right)$
$=$ $O_{p}\left(\left(\frac{\log L}{Lh^{2j-1}}\right)^{1/2}\left(\frac{\log L}{Lh^{2}}\right)^{1/2}\right)$
$=$ $o_{p}\left(\left(\frac{\log L}{Lh^{2j-1}}\right)^{1/2}\right),$
[…]
Proof of Lemma 3.
Given the linearization:
$\widehat{B}^{-1}\left(t|x\right)-B^{-1}\left(t|x\right)=-\frac{\widehat{B}\left(B^{-1}\left(t|x\right)|x\right)-B\left(B^{-1}\left(t|x\right)|x\right)}{B^{\left(1\right)}\left(B^{-1}\left(t|x\right)|x\right)}+o_{p}\left(\sup_{\left(\alpha,x\right)\in\left[0,1\right]\times\mathcal{X}}\left|\widehat{B}\left(\alpha|x\right)-B\left(\alpha|x\right)\right|\right),$
[…]
C. Asymptotic Theory for Risk Aversion Estimator
C.1 Preliminaries
The following class of functions feature in our objective function and is central to the analysis of our semiparametric estimator:
$\mathcal{Q}=\left\{\begin{array}[]{c}q_{\theta,\psi}\left(w,r,x\right)=U_{\theta}\left(w\right)+U_{\theta}^{\left(1\right)}\left(r\right)\psi_{1}\left(\psi_{2}\left(r,x\right),x\right)\left(1-\psi_{2}\left(r,x\right)\right)-U_{\theta}\left(r\right)\text{ }\\ \text{for }\left(\theta,\psi\right)\in\Theta\times\mathcal{D}\end{array}\right\},$ (24)
where $\theta\in\Theta$ is a compact subset of $\mathbb{R}\$and $\psi=\left(\psi_{1},\psi_{2}\right)\in\mathcal{D}=\mathcal{D}_{1}\times\mathcal{D}_{2}$ where $\mathcal{D}_{1}\mathcal{\ }$and $\mathcal{D}_{2}$ respectively abbreviate $\mathcal{D}_{1}\left(\mathcal{A_{\delta}}\times\mathcal{X}\right)\times\mathcal{D}_{2}\left(\mathcal{V_{\delta}}\times\mathcal{X}\right)$.
Before proceeding further, it will be useful to state some conventions on notations. First, to simplify, we will frequently suppress the pointwise arguments of the functions unless its serves to clarify matters. Second, since composite functions are a prominent feature in $\mathcal{Q}$, to facilitate the readers, we use $\circ$ to indicate it in our proofs. Third, given the class of functions we consider are uniformly bounded, we will generically use $\left\|\cdot\right\|_{\infty}$ to denote the sup-norm over all arguments of a real value function over its support. For example, $\sup_{\left(r,x\right)\in\mathcal{V}_{\delta}\times\mathcal{X}}\left|\psi_{1}\left(\psi_{2}\left(r,x\right),x\right)\right|$ will be abbrebiated to $\left\|\psi_{1}\circ\psi_{2}\right\|_{\infty}$. And, to emphasize we are performing calculations with respect to the parameters when applying empirical process theory, we highlight the parameters by making them the arguments of the functions, e.g., $q_{\theta,\psi}\left(\cdot\right)$ becomes $q\left(\theta,\psi\right)$. We will use $\lesssim$ to denote the less than or equal to relation up to a universal constant.
We will show that $\mathcal{Q}$ and related class of functions are $P_{Z}$-Glivenko-Cantelli under S1 and $P_{Z}$-Donsker under S2. We focus our proofs on the Donsker case, which is the more difficult of the two. Let $N_{[]}\left(\varepsilon,\mathcal{Q},L_{2}\left(P_{Z}\right)\right)$ denote the covering numbers with bracketing with respect to the $L_{2}$-norm (see Definition 2.1.6 in VW23). We define the covering numbers for other class of functions using a similar notation. It will be convenient to use $Z_{0}$ to denote a generic subset from $\left(W,R,X\right)$ and $P_{Z_{0}}$ to denote the probability distribution of $Z_{0}$. We will abbreviate $L_{2}\left(P_{Z_{0}}\right)$ to $L_{2}$ when there is no risk of ambiguity. Correspondingly, we use $\left\|\cdot\right\|_{L_{2}}$ to denote the $L_{2}$-norm.
We will prove the Donsker property for a class of functions by showing it has an $L_{2}$-integrable envelope and satisfies the following integral bracketing entropy condition:
$\int_{0}^{\infty}\sqrt{\log N_{[]}\left(\varepsilon,\mathcal{Q},L_{2}\right)}d\varepsilon<\infty,$
see Theorem 2.5.2 in VW23. GC property holds under a simpler to verify requirement that $N_{[]}\left(\varepsilon,\mathcal{Q},L_{1}\right)<\infty$ for all $\varepsilon>0$, see Theorem 2.4.1 in VW23.
The main bracketing number calculations can be traced to the functions in $\mathcal{D}_{j}$, which are assumed to have at least two continuous derivatives under Assumption S2(ii). We denote the class of functions that have $2$ continuous derivatives on $\mathcal{A}_{\delta}$ with an $L_{2}$ integrable envelope by $\mathcal{C}^{2}$.
In what follows, we define $c_{\mathcal{Q}^{0}}$ to be an envelope for functions in $\mathcal{Q}$:
$\left|q_{\theta,\psi}\left(w,r,x\right)\right|$ $\leq$ $c_{\mathcal{Q}^{0}}\left(w,r,x\right),\text{ where}$ (25)
$c_{\mathcal{Q}^{0}}\left(w,r,x\right)$ $=$ $\left|c_{\Theta^{0}}\left(w\right)\right|+\left|c_{\Theta^{0}}\left(r\right)\right|+C_{\mathcal{D}_{1}^{0}}\left|c_{\Theta^{1}}\left(r\right)\right|,$
such that $c_{\Theta^{j}}\left(w\right)=\sup_{\theta\in\Theta}\left|U_{\theta}^{\left(j\right)}\left(w\right)\right|$ and $\sup_{\psi_{i}\in\mathcal{D}_{i}^{j}}\left\|\psi_{i}^{\left(j\right)}\right\|_{\infty}<C_{\mathcal{D}_{i}^{j}}$ for some $C_{\mathcal{D}_{i}^{j}}\$fori=1,2\ i=1,2 and $j=0,1,2$. Under Assumptions S1(iii)-(iv) and S2(i)-(ii), $E\left[c_{\Theta^{j}}\left(W,R,X\right)^{i}\right]<\infty$ and $C_{\mathcal{D}_{i}^{j}}<\infty$ for $i=1,2$ and $j=0,1,2$.
C.2 Additional Lemmas
Lemma C.1. $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ defined in Assumption S2(ii) are $P_{Z}$-Donsker.
Proof of Lemma C.1.
We start by showing the class of functions that $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ are derived from are Donsker. Let $\mathcal{M}_{j}=\{$ $x_{j}\mu\left(\alpha\right)$ for $x_{j}\in\mathcal{X}_{j}$ and $\mu\in\mathcal{C}^{2}$ }\} where $\mathcal{X}_{j}\subseteq\mathbb{R}$ is the support of the $j$-th component in the random vector $X$. It is well known that $\mathcal{C}^{2}$ is a Donsker class of functions. $\mathcal{M}_{j}$ is a class of $\mathcal{C}^{2}$ functions with a bounded fixed multiplier due to compactness of $\mathcal{X}_{j}$. Let $\overline{\mu}$ be an envelope on $\mathcal{C}^{2}$ with $P_{Z_{0}}\overline{\mu}^{2}<\infty$, then $C_{\mathcal{X}_{j}}\overline{\mu}$ is an envelope for $\mathcal{M}_{j}$ where $C_{\mathcal{X}_{j}}=\max_{x_{j}\in\mathcal{X}_{j}}\left|x_{j}\right|$. It follows that $\mathcal{M}_{j}$ is Donsker since brackets $\left\{\left[l_{\jmath},u_{\jmath}\right]\right\}_{\jmath=1}^{N}$ that cover $\mathcal{C}^{2}$ leads to $\left\{\left[x_{j}l_{\jmath},x_{j}u_{\jmath}\right]\right\}_{\jmath=1}^{N}$ which covers $\mathcal{M}_{j}$. Then, $N_{[]}\left(\varepsilon,x_{j}\mathcal{C}_{2},L_{2}\left(P_{Z_{0}}\right)\right)\leq N_{[]}\left(\varepsilon/C_{\mathcal{X}_{j}},\mathcal{C}_{2},L_{2}\left(P_{Z_{0}}\right)\right)$, so the bracketing entropy condition is satisfied since $\log N_{[]}\left(\varepsilon/C_{\mathcal{X}_{j}},\mathcal{C}_{2},L_{2}\left(P_{Z_{0}}\right)\right)\lesssim\varepsilon^{-1}$ (e.g., see Corollary 2.7.2 in VW23). Let $\mathcal{M}=\{$ $m_{0}\left(\alpha\right)+\sum\limits_{j=1}^{D}m_{j}\left(\alpha,x_{j}\right)$ for $x\in\mathcal{X}$ such that $m_{0}\in\mathcal{C}^{2}\left(\mathcal{A}_{\delta}\right)$ and $m_{j}\in\mathcal{M}_{j}$ }\}. $\left(1+\sum\limits_{j=1}^{D}C_{\mathcal{X}_{j}}\right)\overline{\mu}$ is an $L_{2}$-integrable envelope for $\mathcal{M}$. $\mathcal{M}$ is also Donsker due to the preservation of Donsker properties of a finite sum of classes of Donsker functions (VW23, Theorem 2.10.8).
Let $\mathcal{M}^{1}=\{$ $\phi^{\left(1\right)}\left(\alpha\right)m\left(\phi\left(\alpha\right),x\right)$ for $m\in\mathcal{M}\ \}$. A function in $\mathcal{M}^{1}$ takes the form $\frac{\partial}{\partial\alpha}\nu\left(\phi\left(\alpha\right),x\right)$ for $\nu\in\mathcal{D}_{0}$ defined in Assumption S1(iv). Thus, $\mathcal{D}_{1}$ is a subset of $\mathcal{M}^{1}$ when Assumption S2(ii) holds, and $\mathcal{D}_{1}$ will be Donsker if $\mathcal{M}^{1}$ is Donsker (VW23, Theorem 2.10.8). $\mathcal{M}^{1}$ is indeed Donsker. This follows as $\mathcal{M}^{1}$ is a class of Donsker functions with a bounded fixed multiplier, $\phi^{\left(1\right)}$ – whose image is compact on $A_{\delta}$. Note that the composition of $\phi$ in the first argument of $m\in\mathcal{M}\$does not affect the bracketing entropy calculation. This is because $\phi$ is strictly increasing and the composition serves as a change of probability measure that defines the $L_{2}$ norm in the bracketing number calculation, and the bracketing number can be chosen to be independent of the probability measure.
Under Assumption S1(iv), $\mathcal{D}_{2}$ is defined as $\left(t,x\right)\mapsto\nu_{x}^{-1}\left(t\right)$ where $\nu_{x}\left(\alpha\right)=\nu\left(\phi\left(\alpha\right),x\right)$ for $\nu\in\mathcal{D}_{0}$. Noting that $\mathcal{D}_{0}$ is a subset of $\mathcal{M}$. By definition of $\mathcal{D}_{0}$, the partial inverse of $\nu_{x}$ exists and we denote it by $\nu_{x}^{-1}:\mathcal{V}_{\delta}\mapsto\mathcal{A}_{\delta}$. Consider the relation $\nu\left(\phi\left(\nu_{x}^{-1}\left(t\right)\right),x\right)=t$. The Implicit Function Theorem gives,
$\frac{\partial}{\partial x}\nu_{x}^{-1}\left(t\right)=\frac{\mu\left(\phi\left(\nu_{x}^{-1}\left(t\right)\right)\right)}{\phi^{\left(1\right)}\left(\nu_{x}^{-1}\left(t\right)\right)x_{1}^{\top}\frac{\partial}{\partial\alpha}\mu\left(\phi\left(\nu_{x}^{-1}\left(t\right)\right)\right)}.$
Under S1(iv), $\sup_{\left(t,x\right)\in\mathcal{V}_{\delta}\times\mathcal{X}}\frac{\partial}{\partial x}\nu_{x}^{-1}\left(t\right)<\infty$ uniformly for $\nu\in\mathcal{D}_{0}$, which implies functions in $\mathcal{D}_{2}$ is uniformly Lipschitz with respect to the argument on $\mathcal{X}$. We denote the corresponding Lipschitz constant by $K_{\mathcal{X}}$. We will use this Lipschitz property to build suitable brackets covering $\mathcal{D}_{2}$ and show the bracketing entropy is integrable.
For any $\varepsilon>0$, take a $\eta$-net, $\left\{x_{\ell}\right\}_{\ell=1}^{N_{\mathcal{X}}}$, of $\mathcal{X}$ such that $\eta=\varepsilon/3K_{\mathcal{X}}$ and $N_{\mathcal{X}}\lesssim\varepsilon^{-D}$. For each $x_{\ell}$, we can find $N_{\ell}\left(\varepsilon\right)$ brackets to cover all functions $t\mapsto\nu_{x_{\ell}}^{-1}\left(t\right)$ that are uniformly bounded and monotone on $\mathcal{V}_{\delta}$, $\left\{\left[l_{\ell\jmath},u_{\ell\jmath}\right]\right\}_{\jmath=1}^{N_{\ell}\left(\varepsilon\right)}$, such that $\left\|u_{\jmath\ell}-l_{\jmath\ell}\right\|_{L_{2}\left(P_{Z_{0}}\right)}\leq\varepsilon/3$ and $\log N_{\ell}\left(\varepsilon\right)\lesssim\varepsilon^{-1}$ (VW23, Theorem 2.7.9). We can then extend the brackets at $\left\{x_{\ell}\right\}_{\ell=1}^{N_{\mathcal{X}}}$ to brackets on $\mathcal{V}_{\delta}\times\mathcal{X}$. To do this, for any $x$ such that $\left\|x-x_{\ell}\right\|\leq\eta$, define $\underline{l}_{\ell\jmath}\left(t,x\right)=l_{\ell\jmath}\left(t\right)-\varepsilon/3$ and $\overline{u}_{\ell\jmath}\left(t,x\right)=u_{\ell\jmath}\left(t\right)+\varepsilon/3$. By combining the Lipschitz continuity (over $x$) together with the brackets for $\nu_{x_{\ell}}^{-1}$, we have $\underline{l}_{\ell\jmath}\left(t,x\right)\leq\nu_{x}^{-1}\left(t\right)\leq\overline{u}_{\ell\jmath}\left(t,x\right)$ for all $t$ and $\left\|x-x_{\ell}\right\|\leq\eta$. Thus, $\left\{\left[\underline{l}_{\ell\jmath},\overline{u}_{\ell\jmath}\right]\right\}_{\jmath=1}^{N_{\ell}\left(\varepsilon\right)}$ are brackets for $\nu_{x}^{-1}$ on the $\eta$-ball around $x_{\ell}$. Since the union of the $\eta$-balls around $\left\{x_{\ell}\right\}_{\ell=1}^{N_{\mathcal{X}}}$ cover $\mathcal{X}$, we have a cover for $\mathcal{D}_{2}$. Note that $\left\|\overline{u}_{\ell\jmath}-\underline{l}_{\ell\jmath}\right\|_{L_{2}\left(P_{Z_{0}}\right)}\leq\left\|u_{\ell\jmath}-l_{\ell\jmath}\right\|_{L_{2}\left(P_{Z_{0}}\right)}+2\varepsilon/3\leq\varepsilon$. Therefore, $N_{[]}\left(\varepsilon,\mathcal{D}_{2},L_{2}\left(P_{Z_{0}}\right)\right)\leq\sum\limits_{\ell=1}^{N_{\mathcal{X}}}N_{\ell}\left(\varepsilon\right)$ and $\log N_{[]}\left(\varepsilon,\mathcal{D}_{2},L_{2}\left(P_{Z_{0}}\right)\right)\lesssim\varepsilon^{-1}+\log\left(\varepsilon^{-1}\right)$, so that $\int_{0}^{1}\sqrt{\log N_{[]}\left(\varepsilon,\mathcal{D}_{2},L_{2}\left(P_{Z_{0}}\right)\right)}d\varepsilon<\infty$$\blacksquare$
Lemma C.2. Under Assumption S2(ii), $\mathcal{Q}$ is $P_{Z}$-Donsker.
Proof of Lemma C.2.
Each function in $\mathcal{Q}$ is a sum of three terms. Two of them belong to a low complexity class of functions, $\left\{U_{\theta}:\text{ }\theta\in\Theta\right\}$, that are Lipschitz over $\Theta$ with an $L_{2}\left(P_{Z_{0}}\right)$ integrable envelope. Such a class of functions satisfies the entropy bracketing condition (VW23, Thorem 2.7.17) and is Donsker. The remaining term is a product of three functions: $\left(r,x\right)\mapsto U_{\theta}^{\left(1\right)}\left(r\right)\psi_{1}\left(\psi_{2}\left(r,x\right),x\right)\left(1-\psi_{2}\left(r,x\right)\right)$ where $\left(\psi_{1},\psi_{2}\right)\in\mathcal{D}_{1}\times\mathcal{D}_{2}$. $\left\{U_{\theta}^{\left(1\right)}:\theta\in\Theta\right\}$ is Donsker as it is also a Lipschitz-in-parameter class of functions with an $L_{2}\left(P_{Z_{0}}\right)$ integrable envelope. We have already shown in Lemma C.1 that $\mathcal{D}_{2}$ is Donsker, so that $\left(r,x\right)\mapsto 1-\psi_{2}\left(r,x\right)$ is Donsker. We will show below that $\mathcal{D}_{1\circ 2}=\{$ $\left(r,x\right)\mapsto\psi_{1}\left(\psi_{2}\left(r,x\right),x\right)$ for $\left(\psi_{1},\psi_{2}\right)\in\mathcal{D}_{1}\times\mathcal{D}_{2}$ }\} is also Donsker. Together, this ensures $\mathcal{Q}$ is Donsker as permanence properties of Donsker classes apply to finite sums and products of Donsker classes of functions.
$\mathcal{D}_{1\circ 2}$ is a class of composition of functions with the outer map belonging to $\mathcal{D}_{1}$. The $L_{2}\left(P_{Z_{0}}\right)$ integrable envelope for $\mathcal{D}_{1}$ then applies for functions in $\mathcal{D}_{1\circ 2}$. Note that the standard permanence Donsker theorem employed previously (VW23, Theorem 2.10.8) assumes the outer map of the composition function is fixed, ao it does not apply to $\mathcal{D}_{1\circ 2}$. We need to verify Donsker permanence holds uniformly over $\mathcal{D}_{1}$ as well. To do this, take any $\psi_{1},\psi_{1}^{\prime}$ in $\mathcal{D}_{1}$ and $\psi_{2}\in\mathcal{D}_{2}$, we have
$\left\|\psi_{1}\circ\psi_{2}-\psi_{1}^{\prime}\circ\psi_{2}\right\|_{L_{2}\left(P_{Z_{0}}\right)}^{2}$ $=$ $\int\left(\psi_{1}\left(\psi_{2}\left(R,X\right),X\right)-\psi_{1}^{\prime}\left(\psi_{2}\left(R,X\right),X\right)\right)^{2}dP_{Z_{0}}$
$=$ $\int\left(\psi_{1}\left(A,X\right)-\psi_{1}^{\prime}\left(A,X\right)\right)^{2}dP_{\widetilde{Z}_{0}\left(\psi_{2}\right)}.$
The second equality above follows from a change of measure that is valid due to $\psi_{2}$ being strictly increasing. Let $\lambda_{R,X}$ be a dominating measure of $P_{Z_{0}}$, defined as the product of a Lebesgue measure on $\mathcal{V}_{\delta}$ and a general measure on $X$ (possibly containing a counting measure if components of $X$ is discrete). We denote $\frac{dP_{Z_{0}}}{d\lambda_{R,X}}$, a Radon–Nikodym derivative, by $p_{R,X}$. Then, $\frac{dP_{\widetilde{Z}_{0}\left(\psi_{2}\right)}}{d\lambda_{A,X}}$ is a map $\left(\alpha,x\right)\mapsto\frac{p_{R,X}\left(\psi_{2}^{-1}\left(\alpha,x\right),x\right)}{\psi_{2}^{\left(1\right)}\left(\psi_{2}^{-1}\left(\alpha,x\right),x\right)}$, where $\lambda_{A,X}$ is a dominating measure defined similarly to $\lambda_{R,X}$ with a measure on $\mathcal{A}_{\delta}$ instead of $\mathcal{V}_{\delta}$. Under conditions of S1(iv), $\frac{dP_{\widetilde{Z}_{0}\left(\psi_{2}\right)}}{d\lambda_{A,X}}$ is (pointwise) bounded above by some constant $C_{P}<\infty$ uniformly over $\mathcal{D}_{2}$. Hence,
$\left\|\psi_{1}\circ\psi_{2}-\psi_{1}^{\prime}\circ\psi_{2}\right\|_{L_{2}\left(P_{Z_{0}}\right)}=\left\|\psi_{1}-\psi_{1}^{\prime}\right\|_{L_{2}\left(P_{\widetilde{Z}_{0}\left(\psi_{2}\right)}\right)}\leq\sqrt{C_{P}}\left\|\psi_{1}-\psi_{1}^{\prime}\right\|_{L_{2}\left(\lambda_{A,X}\right)}.$
Since $\mathcal{D}_{1\circ 2}$ consists of composition functions with functions from $\mathcal{D}_{1}$ as the outer map, the brackets that cover $\mathcal{D}_{1}$ can be used to cover $\mathcal{D}_{1\circ 2}$. Subsequently, the inequality above implies:
$N_{[]}\left(\varepsilon,\mathcal{D}_{1\circ 2},L_{2}\left(P_{Z_{0}}\right)\right)\leq N_{[]}\left(\varepsilon/\sqrt{C_{P}},\mathcal{D}_{1},L_{2}\left(\lambda_{A,X}\right)\right).$
The bound on the bracketing numbers used to establish Donsker property of $\mathcal{D}_{1}$ in Lemma C.1 applies to any $L_{2}$ norm defined with finite measure. Thus, we have $\int_{0}^{1}\sqrt{\log N_{[]}\left(\varepsilon,\mathcal{D}_{1},L_{2}\left(\lambda_{A,X}\right)\right)}d\varepsilon<\infty$$\blacksquare$
Lemma C.3. Under Assumption S1(iv), $\mathcal{Q}$ is $P_{Z}$-Glivenko-Cantelli.
Proof of Lemma C.3.
[…]
C.3 Proofs of results
Proof of Lemma 4.
Since $\Theta$ is compact and $\theta\mapsto Q\left(\theta,\psi_{0}\right)$ is continuous, it suffices to show that $Q\left(\theta_{0},\psi_{0}\right)\leq Q\left(\theta,\psi_{0}\right)$ and the equality holds if and only if $\theta=\theta_{0}$. For this, as done in the proof of Proposition 2, let us first omit $X$ and consider the parameterized version of (15),
$h_{\theta}(r,w)=U_{\theta}(w)+U_{\theta}^{\left(1\right)}(r)\frac{1-F(r)}{f(r)}-U_{\theta}(r).$
By Proposition 2, for any $w$, suppose $r$ is the corresponding optimal reserve price when the seller’s risk aversion is $\theta_{0}$, there must be
$h_{\theta_{0}}(r,w)=0.$
We first show that $h_{\theta}(r,w)\neq 0$ whenever $\theta\neq\theta_{0}$. Consider $\theta>\theta_{0}$, there exists a function $\zeta(\cdot)$ satisfying condition (a) of S1(iii) su that $U_{\theta}(\cdot)=\zeta(U_{\theta_{0}}(\cdot))$. Therefore, we have
$U_{\theta}(r)-U_{\theta}^{(1)}(r)\frac{1-F(r)}{f(r)}$ $=\zeta(U_{\theta_{0}}(r))-\zeta^{(1)}(U_{\theta_{0}}(r))U_{\theta_{0}}^{(1)}(r)\frac{1-F(r)}{f(r)}$
$>\zeta\left(U_{\theta_{0}}(r)-U_{\theta_{0}}^{(1)}(r)\frac{1-F(r)}{f(r)}\right)$
$=\zeta(U_{\theta_{0}}(w))$
$=U_{\theta}(w),$
which implies
$h_{\theta}(r,w)=U_{\theta}(w)+U_{\theta}^{\left(1\right)}(r)\frac{1-F(r)}{f(r)}-U_{\theta}(r)<0.$
A similar argument can be applied to the case $\theta<\theta_{0}$, so that
$U_{\theta_{0}}(w)$ $=U_{\theta_{0}}(r)-U_{\theta_{0}}^{(1)}(r)\frac{1-F(r)}{f(r)}$
$=\zeta(U_{\theta}(r))-\zeta^{(1)}(U_{\theta}(r))U_{\theta}^{(1)}(r)\frac{1-F(r)}{f(r)}$
$>\zeta\left(U_{\theta}(r)-U_{\theta}^{(1)}(r)\frac{1-F(r)}{f(r)}\right).$
and we have,
$U_{\theta}(r)-U_{\theta}^{\left(1\right)}(r)\frac{1-F(r)}{f(r)}<\zeta^{-1}(U_{\theta_{0}}(w))=U_{\theta}(w),$
which implies
$h_{\theta}(r,w)=U_{\theta}(w)+U_{\theta}^{\left(1\right)}(r)\frac{1-F(r)}{f(r)}-U_{\theta}(r)>0.$
Inputting back the dependence on $X$, and noting that $\frac{1-F(r|x)}{f(r|x)}=\psi_{10}\left(\psi_{20}\left(r,x\right),x\right)\left(1-\psi_{20}\left(r,x\right)\right)$, the argument above implies that $q\left(z,\theta,\psi_{0}\right)\neq 0$ when $\theta\neq\theta_{0}$ for $z\in\mathcal{Z}$, which means $Q\left(\theta,\psi_{0}\right)>0$ when $\theta\neq\theta_{0}$$\blacksquare$
Proof of Lemma 5.
From the triangle inequality, we have
$\left|Q_{L}\left(\theta,\widehat{\psi}\right)-Q\left(\theta,\psi_{0}\right)\right|\leq\left|Q_{L}\left(\theta,\widehat{\psi}\right)-Q\left(\theta,\widehat{\psi}\right)\right|+\left|Q\left(\theta,\widehat{\psi}\right)-Q\left(\theta,\psi_{0}\right)\right|.$
Under the assumed bandwidth condition, $\widehat{\psi}_{i}\in\mathcal{D}_{i}$ w.p.a. $1$ for $i=1,2$. In particular, as discussed in Appendix C.1 of GG22, AQR estimators are smooth w.p.a. $1$, which ensures,
$\left|Q_{L}\left(\theta,\widehat{\psi}\right)-Q\left(\theta,\widehat{\psi}\right)\right|\leq\sup_{\theta\in\Theta,\psi\in\mathcal{D}_{1}\times\mathcal{D}_{2}}\left|Q_{L}\left(\theta,\psi\right)-Q\left(\theta,\psi\right)\right|=o_{p}\left(1\right).$
The uniform convergence is a consequence of Lemma C.3.
[…]
$\left|q\left(\theta,\widehat{\psi}\right)^{2}-q\left(\theta,\psi_{0}\right)^{2}\right|\leq 2c_{\mathcal{Q}^{0}}\left|q\left(\theta,\widehat{\psi}\right)-q\left(\theta,\psi_{0}\right)\right|,$
due to the triangle inequality and bounding by an envelope of $\mathcal{Q}$. It will be instructive here to prove that $q$ is (pointwise) Lipschitz continuous with respect to $\psi$ for any $\theta$ in the following sense:
$\left|q\left(\theta,\psi\right)-q\left(\theta,\psi^{\prime}\right)\right|\leq c\left[\left\|\psi_{1}-\psi_{1}^{\prime}\right\|_{\infty}+\left\|\psi_{2}-\psi_{2}^{\prime}\right\|_{\infty}\right],$
for some function $c$ such that $E\left[c^{2}\right]<\infty$. The main calculation involves showing that $T\left(\psi\right)=\left(\psi_{1}\circ\psi_{2}\right)\left(1-\psi_{2}\right)$ is Lipschitz continuous, which will also be useful in other parts of the appendix, so we provide some details here.
Take any $\psi,\psi^{\prime}$ and consider,
$T\left(\psi\right)-T\left(\psi^{\prime}\right)$ $=$ $\left(\psi_{1}\circ\psi_{2}-\psi_{1}^{\prime}\circ\psi_{2}\right)\left(1-\psi_{2}\right)$
$+\left(\psi_{1}^{\prime}\circ\psi_{2}-\psi_{1}^{\prime}\circ\psi_{2}^{\prime}\right)\left(1-\psi_{2}\right)$
$+\left(\psi_{1}^{\prime}\circ\psi_{2}^{\prime}\right)\left(\left(1-\psi_{2}\right)-\left(1-\psi_{2}^{\prime}\right)\right).$
Note that the image of functions in $\mathcal{D}_{2}$ is in $A_{\delta}$ so that $\max\{\sup_{\psi_{2}\in\mathcal{D}_{2}}\left\|1-\psi_{2}\right\|_{\infty},\sup_{\psi_{2}\in\mathcal{D}_{2}}\left\|\psi_{2}\right\|_{\infty}\}<1$. The first term of the sum above then satisfies $\left\|\left(\psi_{1}\circ\psi_{2}-\psi_{1}^{\prime}\circ\psi_{2}\right)\left(1-\psi_{2}\right)\right\|_{{}_{\infty}}\leq\left\|\psi_{1}-\psi_{1}^{\prime}\right\|_{\infty}$. For the second term, we can apply a mean value expansion and impose the upper bound on $\psi_{1}^{\left(1\right)}$ to obtain $\left\|\psi_{1}^{\prime}\circ\psi_{2}-\psi_{1}^{\prime}\circ\psi_{2}^{\prime}\right\|_{{}_{\infty}}\leq C_{\mathcal{D}_{1}^{1}}\left\|\psi_{2}-\psi_{2}^{\prime}\right\|_{\infty}$. The last term is bounded above by $C_{\mathcal{D}_{1}^{1}}\left\|\psi_{2}-\psi_{2}^{\prime}\right\|_{\infty}$, since $\psi_{1}^{\prime}$ is bounded above by $C_{\mathcal{D}_{1}^{0}}$. Together, they imply
$\left\|T\left(\psi\right)-T\left(\psi^{\prime}\right)\right\|_{\infty}\leq\left(1+C_{\mathcal{D}_{1}^{1}}+C_{\mathcal{D}_{1}^{0}}\right)\left(\left\|\psi_{1}-\psi_{1}^{\prime}\right\|_{\infty}+\left\|\psi_{2}-\psi_{2}^{\prime}\right\|_{\infty}\right).$ (26)
By S1(iii), $\left|U_{\theta}^{\left(1\right)}\right|\leq c_{\Theta^{1}}$. We then have,
$\left|q\left(\theta,\psi\right)-q\left(\theta,\psi^{\prime}\right)\right|\leq\left(1+C_{\mathcal{D}_{1}^{1}}+C_{\mathcal{D}_{1}^{0}}\right)c_{\Theta^{1}}\left[\left\|\psi_{1}-\psi_{1}^{\prime}\right\|_{\infty}+\left\|\psi_{2}-\psi_{2}^{\prime}\right\|_{\infty}\right],$ (27)
as desired.
Since $\widehat{\psi}_{2}\in\mathcal{D}_{2}$ w.p.a. $1$, it follows that,
$\sup_{\theta\in\Theta}\left|Q\left(\theta,\widehat{\psi}\right)-Q\left(\theta,\psi_{0}\right)\right|$ $\leq$ $2\left(1+C_{\mathcal{D}_{1}^{1}}+C_{\mathcal{D}_{1}^{0}}\right)E\left[c_{\mathcal{Q}^{0}}^{2}\right]^{1/2}E\left[c_{\Theta^{1}}^{2}\right]^{1/2}$
$\times\left(E\left[\left\|\widehat{\psi}_{1}-\psi_{10}\right\|_{\infty}\right]+E\left[\left\|\widehat{\psi}_{2}-\psi_{20}\right\|_{\infty}\right]\right)$
$=$ $o\left(1\right),$
[…]
Proof of Theorem 1.
[…]
Proof of Lemma 6.
Let us define $Y_{L}\left(\theta,\psi\right)=\frac{1}{L}\sum\limits_{l=1}^{L}y\left(Z_{l},\theta,\psi\right)$, where $y\left(Z_{l},\theta,\psi\right)=q\left(Z_{l},\theta,\psi\right)\frac{\partial}{\partial\theta}q\left(Z_{l},\theta,\psi\right)$, and let $Y\left(\theta,\psi\right)=\int q\left(Z_{l},\theta,\psi\right)\frac{\partial}{\partial\theta}q\left(Z_{l},\theta,\psi\right)dP_{Z}$. We want to show $Y\left(\theta_{0},\widehat{\psi}\right)=\dot{Y}\left(\theta_{0},\psi_{0}\right)\left[\widehat{\psi}-\psi_{0}\right]+o_{p}\left(L^{-1/2}\right)$, where $\dot{Y}\left(\theta_{0},\psi_{0}\right)\left[\widehat{\psi}-\psi_{0}\right]$ denotes the pathwise derivative of $Y$ at $\left(\theta_{0},\psi_{0}\right)$ in direction $\left[\widehat{\psi}-\psi_{0}\right]$, and $\sqrt{L}\dot{Y}\left(\theta_{0},\psi_{0}\right)\left[\widehat{\psi}-\psi_{0}\right]$ satisfies a suitable CLT.
To do this, we need to derive $\dot{Y}\left(\theta_{0},\psi_{0}\right)\left[\widehat{\psi}-\psi_{0}\right]$. The main functional components in $Y$ can be traced to $T\left(\psi\right)=\left(\psi_{1}\circ\psi_{2}\right)\left(1-\psi_{2}\right)$, which appears in both $q$ and $\frac{\partial}{\partial\theta}q$. Let us first calculate the pathwise derivative of $T$ at $\psi$ in direction $\omega$. It follows from simple algebra that
$T\left(\psi+\epsilon\omega\right)-T\left(\psi\right)=\epsilon\omega_{1}\circ\psi_{2}\left(1-\psi_{2}\right)+\epsilon\left(\psi_{1}^{\left(1\right)}\circ\psi_{2}\left(1-\psi_{2}\right)-\psi_{1}\circ\psi_{2}\right)\omega_{2}+O\left(\epsilon^{2}\right),$
so that $\lim_{\epsilon\rightarrow 0}\frac{T\left(\psi+\epsilon\omega\right)-T\left(\psi\right)}{\epsilon}$ exists and is equal to
$\dot{T}\left(\psi\right)\left[\omega\right]=\omega_{1}\circ\psi_{2}\left(1-\psi_{2}\right)+\left(\psi_{1}^{\left(1\right)}\circ\psi_{2}\left(1-\psi_{2}\right)-\psi_{1}\circ\psi_{2}\right)\omega_{2}.$
The perturbation $T\left(\psi+\epsilon\omega\right)-T\left(\psi\right)$ can also be refined to obtain the second derivative. It can be checked through some more algebra that the $O\left(\epsilon^{2}\right)$ term in the display above consists of $\epsilon^{2}\Xi\left(\omega\right)+O\left(\epsilon^{3}\right)$, where
$\Xi\left(\omega\right)=\left(\omega_{1}^{\left(1\right)}\circ\psi_{2}\left(1-\psi_{2}\right)-\omega_{1}\circ\psi_{2}\right)\omega_{2}+\left(\frac{1}{2}\psi_{1}^{\left(2\right)}\circ\psi_{2}\left(1-\psi_{2}\right)-\psi_{1}^{\left(1\right)}\circ\psi_{2}\right)\omega_{2}^{2}.$
Focusing on the derivative at $\psi_{0}$, when $\widehat{\psi}-\psi_{0}=\epsilon\omega$, we have
$\Xi\left(\epsilon\omega\right)$ $=$ $O_{p}\left(\left\|\epsilon\omega_{1}\times\epsilon\omega_{2}\right\|_{\infty}+\left\|\epsilon\omega_{1}^{\left(1\right)}\times\epsilon\omega_{2}\right\|_{\infty}+\left\|\left(\epsilon\omega_{2}\right)^{2}\right\|_{\infty}\right)$
$=$ $O_{p}\left(\left\|\left(\widehat{\psi}_{1}^{\left(1\right)}-\psi_{10}^{\left(1\right)}\right)\times\left(\widehat{\psi}_{2}-\psi_{20}\right)\right\|_{\infty}\right),$
where $\left\|\epsilon\omega_{1}^{\left(1\right)}\times\epsilon\omega_{2}\right\|_{\infty}$ is the leading term, as it can be shown that $\left\|\widehat{\psi}_{1}^{\left(1\right)}-\psi_{10}^{\left(1\right)}\right\|_{\infty}=O_{p}\left(\sqrt{\frac{\log L}{Lh^{3}}}+h^{s-1}\right)$ and $\left\|\widehat{\psi}_{2}-\psi_{20}\right\|_{\infty}=O_{p}\left(\sqrt{\frac{\log L}{L}}+h^{s+1}\right)$; these rates correspond to the convergence rates in Proposition 4, with $j=2$, and Proposition 5 respectively. We then have $\Xi\left(\epsilon\omega\right)=O_{p}\left(\frac{\log L}{Lh^{3/2}}+h^{2s}\right)=o_{p}\left(L^{-1/2}\right)$, where the latter holds due to our assumption on $h$. This ensures that $T\left(\widehat{\psi}\right)-T\left(\psi_{0}\right)=\dot{T}\left(\psi_{0}\right)\left[\widehat{\psi}-\psi_{0}\right]+o_{p}\left(L^{-1/2}\right)$.
With $\dot{T}\left(\psi\right)\left[\omega\right]$ on hand, it follows immediately that the pathwise derivatives of $q\left(\theta,\cdot\right)$ and $\frac{\partial}{\partial\theta}q\left(\theta,\cdot\right)$ at $\psi$ in direction $\omega$ are $U_{\theta}^{\left(1\right)}\dot{T}\left(\psi\right)\left[\omega\right]$ and $U_{\theta}^{\left(2\right)}\dot{T}\left(\psi\right)\left[\omega\right]$ respectively for any $\theta$. The pathwise derivatives of $Y\left(\theta,\cdot\right)$ can then be readily obtained from applying the product rule of differentiation to $q\left(\theta,\cdot\right)\frac{\partial}{\partial\theta}q\left(\theta,\cdot\right)$ and integrating this over $P_{Z}$. Specifically,
$\lim_{\epsilon\rightarrow 0}\frac{Y\left(\theta,\psi+\epsilon\omega\right)-Y\left(\theta,\psi\right)}{\epsilon}=\dot{Y}\left(\theta,\psi\right)\left[\omega\right]=\int\frac{\partial}{\partial\theta}q\left(\theta,\psi\right)U_{\theta}^{\left(1\right)}\dot{T}\left(\psi\right)\left[\omega\right]+q\left(\theta,\psi\right)U_{\theta}^{\left(2\right)}\dot{T}\left(\psi\right)\left[\omega\right]dP_{Z}.$
Focusing on the derivative of $Y$ at $\psi_{0}$ in direction $\left[\widehat{\psi}-\psi_{0}\right]$ when $\theta=\theta_{0}$, this yields:
$\dot{Y}\left(\theta_{0},\psi_{0}\right)\left[\widehat{\psi}-\psi_{0}\right]$ $=$ $\int c_{Y_{1}}\left(\widehat{\psi}_{1}-\psi_{10}\right)\circ\psi_{20}+c_{Y_{2}}\left(\widehat{\psi}_{2}-\psi_{20}\right)dP_{Z},\text{ where}$
$c_{Y_{1}}$ $=$ $\left(1-\psi_{20}\right)\left(\frac{\partial}{\partial\theta}q\left(\theta_{0},\psi_{0}\right)U_{\theta_{0}}^{\left(1\right)}+q\left(\theta_{0},\psi_{0}\right)U_{\theta_{0}}^{\left(2\right)}\right),$ (28)
$c_{Y_{2}}$ $=$ $\left(\psi_{10}^{\left(1\right)}\circ\psi_{20}\left(1-\psi_{20}\right)-\psi_{10}\circ\psi_{20}\right)\left(\frac{\partial}{\partial\theta}q\left(\theta_{0},\psi_{0}\right)U_{\theta_{0}}^{\left(1\right)}+q\left(\theta_{0},\psi_{0}\right)U_{\theta_{0}}^{\left(2\right)}\right).$ (29)
Moreover, $Y\left(\theta_{0},\widehat{\psi}\right)-Y\left(\theta_{0},\psi_{0}\right)=\dot{T}\left(\psi_{0}\right)\left[\widehat{\psi}-\psi_{0}\right]+o_{p}\left(L^{-1/2}\right)$.
To show asymptotic normality, we apply Theorem 4 of GG22 where they prove a general CLT holds for an integral functional of the bids quantile and its derivative. To make clear why their result applies, it shall be useful to input the pointwise arguments of our pathwise derivative,
$\dot{Y}\left(\theta_{0},\psi_{0}\right)\left[\widehat{\psi}-\psi_{0}\right]=\int\left.\begin{array}[]{c}c_{Y_{1}}\left(w,r,x\right)\left(\widehat{\psi}_{1}\left(\psi_{20}\left(r,x\right),x\right)-\psi_{10}\left(\psi_{20}\left(r,x\right),x\right)\right)\\ +c_{Y_{2}}\left(w,r,x\right)\left(\widehat{\psi}_{2}\left(r,x\right)-\psi_{20}\left(r,x\right)\right)\end{array}\right.P_{R,W,X}\left(dr,dw,dx\right).$
Next, we re-write the integrand above as a functional of the observed bid’s quantile and its derivative,B\ B and $B^{\left(1\right)}$ respectively. Recalling the relations between bid and valuation quantiles, see (10) and (11), let us define
$\Gamma\left(w,r,x,B,B^{\left(1\right)}\right)=c_{Y_{1}}\left(w,r,x\right)\phi^{\prime}\left(\psi_{20}\left(r,x\right)\right)B^{\left(1\right)}\left(\psi_{20}\left(r,x\right)|x\right)+c_{Y_{2}}\left(w,r,x\right)\phi^{-1}(B^{-1}\left(r,x\right)).$
Our $\Gamma$ is analogous to the integrand that defines the functional in equation (3.1) of GG22. It is easy to see that $\Gamma$ is twice continuously differentiable with respect to $B,B^{\left(1\right)}$, so that $\Gamma$ satisfies Assumption F in their article. To apply their result, we define the following functionals:
$\widehat{\xi}\left(w,x\right)$ $=$ $\int\Gamma\left(w,r,x,\widehat{B},\widehat{B}^{\left(1\right)}\right)P_{R|W,X}\left(dr|w,x\right),$
$\xi_{0}\left(w,x\right)$ $=$ $\int\Gamma\left(w,r,x,B,B^{\left(1\right)}\right)P_{R|W,X}\left(dr|w,x\right),$
$\widehat{\xi}$ $=$ $\int\widehat{\xi}\left(w,x\right)P_{W,X}\left(dw,dx\right),$
$\xi_{0}$ $=$ $\int\xi_{0}\left(w,x\right)P_{W,X}\left(dw,dx\right).\text{ \ }$
[…]
Proof of Theorem 2.
We use $Y_{L},Y,y$ as defined in the proof of Lemma 6. We begin with a statement that $\mathcal{Q}_{1}=\left\{\frac{\partial}{\partial\theta}q\text{ for }q\in\mathcal{Q}\right\}$ is a Donsker class, which subsequently implies $\mathcal{Q}_{2}=\left\{q\frac{\partial}{\partial\theta}q\text{ for }q\in\mathcal{Q}\right\}$ is Donsker due to preservation of the Donsker properties of product of two Donsker classes of functions (VW23, Theorem 2.10.8). We omit details on proving $\mathcal{Q}_{1}$ is Donsker, as the argument used in proving Lemma C.2 directly applies once $U_{\theta}^{\left(j\right)}$ is replaced by $U_{\theta}^{\left(j+1\right)}$.
[…]
$0=Y_{L}\left(\widehat{\theta},\widehat{\psi}\right)=Y\left(\widehat{\theta},\widehat{\psi}\right)+o_{p}\left(1/\sqrt{L}\right),$ (30)
since $\left|\widehat{\theta}-\theta_{0}\right|=o_{p}\left(1\right)$ and $\left\|\widehat{\psi}_{i}-\psi_{i0}\right\|_{\infty}=o_{p}\left(1\right)$ for $i=1,2$. This is valid as long as we can show $\left\|y-y^{\prime}\right\|_{L_{2}\left(P_{Z}\right)}\lesssim\rho\left(\left(\theta,\psi\right),\left(\theta^{\prime},\psi^{\prime}\right)\right)$ in a neighborhood of $\left(\theta_{0},\psi_{0}\right)$ for some semi-metric $\rho$. We use $\rho\left(\left(\theta,\psi\right),\left(\theta^{\prime},\psi^{\prime}\right)\right)=\left|\theta-\theta^{\prime}\right|+\left\|\psi_{1}-\psi_{1}^{\prime}\right\|_{\infty}+\left\|\psi_{2}-\psi_{2}^{\prime}\right\|_{\infty}$, hence it suffices to show $\left(\theta,\psi\right)\mapsto q\left(\theta,\psi\right)\frac{\partial}{\partial\theta}q\left(\theta,\psi\right)$ is Lipschitz with respect to $\rho$. To do this, it is enough to verify that $q$ and $\frac{\partial}{\partial\theta}q$ are Lipschitz, because it readily follows that a product of bounded Lipschitz functions is Lipschitz.
In the proof of Lemma 5, using (26), we have shown in (27) that $q$ is Lipschitz in $\psi$ for fixed $\theta$. It is easy to extend this over $\Theta$, as $U_{\theta}^{\left(j\right)}$ is Lipschitz due to continuity of $U_{\theta}^{\left(j+1\right)}$ on $\Theta$ for $j=0,1$. It is easy to verify that,
$\left|q\left(\theta,\psi\right)-q\left(\theta^{\prime},\psi^{\prime}\right)\right|$ $\leq$ $\left(2c_{\Theta^{1}}+C_{\mathcal{D}_{1}^{0}}c_{\Theta^{2}}\right)\left|\theta-\theta^{\prime}\right|$
$+\left(1+C_{\mathcal{D}_{1}^{1}}+C_{\mathcal{D}_{1}^{0}}\right)c_{\Theta^{1}}\left[\left\|\psi_{1}-\psi_{1}^{\prime}\right\|_{\infty}+\left\|\psi_{2}-\psi_{2}^{\prime}\right\|_{\infty}\right].$
Since $c_{\Theta^{j}}$ are uniformly bounded functions under S1(iii), we have
$\left|q\left(\theta,\psi\right)-q\left(\theta^{\prime},\psi^{\prime}\right)\right|\lesssim\rho\left(\left(\theta,\psi\right),\left(\theta^{\prime},\psi^{\prime}\right)\right),$
as desired. Proving that $\frac{\partial}{\partial\theta}q$ is Lipschitz can be done analogously. The only difference is $U_{\theta}^{\left(j\right)}$ in $q$ becomes $U_{\theta}^{\left(j+1\right)}$ in $\frac{\partial}{\partial\theta}q$, which gives
$\left|\frac{\partial}{\partial\theta}q\left(\theta,\psi\right)-\frac{\partial}{\partial\theta}q\left(\theta^{\prime},\psi^{\prime}\right)\right|$ $\leq$ $\left(2c_{\Theta^{2}}+C_{\mathcal{D}_{1}^{0}}c_{\Theta^{3}}\right)\left|\theta-\theta^{\prime}\right|$
$+\left(1+C_{\mathcal{D}_{1}^{1}}+C_{\mathcal{D}_{1}^{0}}\right)c_{\Theta^{2}}\left[\left\|\psi_{1}-\psi_{1}^{\prime}\right\|_{\infty}+\left\|\psi_{2}-\psi_{2}^{\prime}\right\|_{\infty}\right],$
where the extra smoothness condition on $U_{\theta}$ is assumed in S2(i). We can then conclude that $\left|y\left(\theta,\psi\right)-y\left(\theta^{\prime},\psi^{\prime}\right)\right|\lesssim\rho\left(\left(\theta,\psi\right),\left(\theta^{\prime},\psi^{\prime}\right)\right)$ and (30) holds.
[…]
$0$ $=$ $Y\left(\theta_{0},\widehat{\psi}\right)+\frac{\partial}{\partial\theta}Y\left(\widetilde{\theta},\widehat{\psi}\right)\left(\widehat{\theta}-\theta_{0}\right)+o_{p}\left(1/\sqrt{L}\right),$
$\widehat{\theta}-\theta_{0}$ $=$ $\left[\frac{\partial}{\partial\theta}Y\left(\widetilde{\theta},\widehat{\psi}\right)\right]^{-1}Y\left(\theta_{0},\widehat{\psi}\right)+o_{p}\left(1/\sqrt{L}\right).$
To complete the proof, we will show that,
$\sqrt{L}\left(\widehat{\theta}-\theta_{0}\right)=\left[\frac{\partial}{\partial\theta}Y\left(\theta_{0},\psi_{0}\right)\right]^{-1}\sqrt{L}\dot{Y}\left(\theta_{0},\psi_{0}\right)\left[\widehat{\psi}-\psi_{0}\right]+o_{p}\left(1/\sqrt{L}\right),$
where $\sqrt{L}\dot{Y}\left(\theta_{0},\psi_{0}\right)\left[\widehat{\psi}-\psi_{0}\right]$ converges in distribution to a normal random variable as established by Lemma 6. I.e., we are left to prove that $\left|\frac{\partial}{\partial\theta}Y\left(\widetilde{\theta},\widehat{\psi}\right)-\frac{\partial}{\partial\theta}Y\left(\theta_{0},\psi_{0}\right)\right|=o_{p}\left(1\right)$ where $\frac{\partial}{\partial\theta}Y\left(\theta_{0},\psi_{0}\right)$ is invertible and the linearization error is $o_{p}\left(L^{-1/2}\right)$.
Adding null to $\frac{\partial}{\partial\theta}Y\left(\widetilde{\theta},\widehat{\psi}\right)-\frac{\partial}{\partial\theta}Y\left(\theta_{0},\psi_{0}\right)$ and apply the triangle inequality gives,
$\left|\frac{\partial}{\partial\theta}Y\left(\widetilde{\theta},\widehat{\psi}\right)-\frac{\partial}{\partial\theta}Y\left(\theta_{0},\psi_{0}\right)\right|\leq\left|\frac{\partial}{\partial\theta}Y\left(\widetilde{\theta},\widehat{\psi}\right)-\frac{\partial}{\partial\theta}Y\left(\widetilde{\theta},\psi_{0}\right)\right|+\left|\frac{\partial}{\partial\theta}Y\left(\widetilde{\theta},\psi_{0}\right)-\frac{\partial}{\partial\theta}Y\left(\theta_{0},\psi_{0}\right)\right|.$
We will prove that $\sup_{\theta\in\Theta}\left|\frac{\partial}{\partial\theta}Y\left(\theta,\widehat{\psi}\right)-\frac{\partial}{\partial\theta}Y\left(\theta,\psi_{0}\right)\right|=o_{p}\left(1\right)$ and $\left|\frac{\partial}{\partial\theta}Y\left(\widetilde{\theta},\psi_{0}\right)-\frac{\partial}{\partial\theta}Y\left(\theta_{0},\psi_{0}\right)\right|=o_{p}\left(1\right)$, which implies $\left|\frac{\partial}{\partial\theta}Y\left(\widetilde{\theta},\widehat{\psi}\right)-\frac{\partial}{\partial\theta}Y\left(\theta_{0},\psi_{0}\right)\right|=o_{p}\left(1\right)$.
First, let us write $\frac{\partial}{\partial\theta}y$ in terms of $q$ and its derivatives:
$\frac{\partial}{\partial\theta}y\left(\theta,\psi\right)=q\left(\theta,\psi\right)\frac{\partial^{2}}{\partial\theta^{2}}q\left(\theta,\psi\right)+\left(\frac{\partial}{\partial\theta}q\left(\theta,\psi\right)\right)^{2}.$
Note that $\frac{\partial}{\partial\theta}y\left(\theta_{0},\psi_{0}\right)=\left(\frac{\partial}{\partial\theta}q\left(\theta_{0},\psi_{0}\right)\right)^{2}$, since $q\left(\theta_{0},\psi_{0}\right)=0$ a.s., and $\int\left(\frac{\partial}{\partial\theta}q\left(\theta_{0},\psi_{0}\right)\right)^{2}dP_{Z}$ is invertible under S2(iii).
Consider any $\theta$, using elementary algebra, we have
$\left|\frac{\partial}{\partial\theta}y\left(\theta,\widehat{\psi}\right)-\frac{\partial}{\partial\theta}y\left(\theta,\psi_{0}\right)\right|$ $\leq$ $c_{\mathcal{Q}^{0}}\left|\frac{\partial^{2}}{\partial\theta^{2}}q\left(\theta,\widehat{\psi}\right)-\frac{\partial^{2}}{\partial\theta^{2}}q\left(\theta,\psi_{0}\right)\right|+c_{\mathcal{Q}^{2}}\left|q\left(\theta,\widehat{\psi}\right)-q\left(\theta_{0},\psi_{0}\right)\right|$
$+2c_{\mathcal{Q}^{1}}\left|\frac{\partial}{\partial\theta}q\left(\theta,\widehat{\psi}\right)-\frac{\partial}{\partial\theta}q\left(\theta_{0},\psi_{0}\right)\right|,\text{ where\ }$
$\sup_{\left(\theta,\psi_{1},\psi_{2}\right)\in\Theta\times\mathcal{D}_{1}\times\mathcal{D}_{2}}\left|\frac{\partial^{j}}{\partial\theta^{j}}q\left(\theta,\psi\right)\right|\leq c_{\mathcal{Q}^{j}}$ for $j=1,2$ such that for all $\left(w,r,x\right)$,
$c_{\mathcal{Q}^{j}}\left(w,r,x\right)=\left|c_{\Theta^{j}}\left(w\right)\right|+\left|c_{\Theta^{j}}\left(r\right)\right|+C_{\mathcal{D}_{1}^{0}}\left|c_{\Theta^{j+1}}\left(r\right)\right|.$
$c_{\mathcal{Q}^{j}}$ extends (25) where $c_{\mathcal{Q}^{0}}$ is defined. The same argument for proving $q\left(\theta,\psi\right)$ is Lipschitz in $\psi$ for all $\theta$ also applies to $\frac{\partial^{j}}{\partial\theta^{j}}q\left(\theta,\psi\right)$, giving us
$\left|\frac{\partial^{j}}{\partial\theta^{j}}q\left(\theta,\psi\right)-\frac{\partial^{j}}{\partial\theta^{j}}q\left(\theta,\psi^{\prime}\right)\right|\leq\left(1+C_{\mathcal{D}_{1}^{1}}+C_{\mathcal{D}_{1}^{0}}\right)c_{\Theta^{j+1}}\left[\left\|\psi_{1}-\psi_{1}^{\prime}\right\|_{\infty}+\left\|\psi_{2}-\psi_{2}^{\prime}\right\|_{\infty}\right],\text{ for }j=1,2,$
which extends (27) in Lemma 5 when $j=0$. Since $\int\left|c_{\Theta^{j+1}}\right|^{2}dP_{Z_{0}}<\infty$, we have $\int\left|c_{\mathcal{Q}^{j}}\right|^{2}dP_{Z_{0}}<\infty$ and $\int\left|c_{\mathcal{Q}^{j}}c_{\Theta^{l+1}}\right|dP_{Z_{0}}<\infty$ by repeated applications of Cauchy Schwarz inequality for $j,l=0,1,2$. Then it follows from Jensen’s inequality that $\sup_{\theta\in\Theta}\left|\frac{\partial}{\partial\theta}Y\left(\theta,\widehat{\psi}\right)-\frac{\partial}{\partial\theta}Y\left(\theta,\psi_{0}\right)\right|=O_{p}(\left\|\widehat{\psi}_{1}-\psi_{10}\right\|_{\infty}+\left\|\widehat{\psi}_{2}-\psi_{20}\right\|_{\infty})=o_{p}\left(1\right)$.
For $\left|\frac{\partial}{\partial\theta}Y\left(\widetilde{\theta},\psi_{0}\right)-\frac{\partial}{\partial\theta}Y\left(\theta_{0},\psi_{0}\right)\right|$, using $y\left(\theta_{0},\psi_{0}\right)=0$ a.s. and a similar argument to previously, we have:
$\left|\frac{\partial}{\partial\theta}y\left(\widetilde{\theta},\psi_{0}\right)-\frac{\partial}{\partial\theta}y\left(\theta_{0},\psi_{0}\right)\right|$ $\leq$ $c_{\mathcal{Q}^{2}}\left|q\left(\widetilde{\theta},\psi_{0}\right)-q\left(\theta_{0},\psi_{0}\right)\right|+2c_{\mathcal{Q}^{1}}\left|\frac{\partial}{\partial\theta}q\left(\widetilde{\theta},\psi_{0}\right)-\frac{\partial}{\partial\theta}q\left(\theta_{0},\psi_{0}\right)\right|$
$\leq$ $3c_{\mathcal{Q}^{1}}c_{\mathcal{Q}^{2}}\left|\widetilde{\theta}-\theta_{0}\right|.$
Integrating the above with respect to $P_{Z}$ and apply Jensen’s inequality gives,
$\left|\frac{\partial}{\partial\theta}Y\left(\widetilde{\theta},\psi_{0}\right)-\frac{\partial}{\partial\theta}Y\left(\theta_{0},\psi_{0}\right)\right|=O_{p}\left(\left|\widetilde{\theta}-\theta_{0}\right|\right),$
where Cauchy Schwarz inequality ensures $\int\left|c_{\mathcal{Q}^{1}}c_{\mathcal{Q}^{2}}\right|dP_{Z}<\infty$. The proof is completed as $\left|\widetilde{\theta}-\theta_{0}\right|=o_{p}\left(1\right)$$\blacksquare$